### v1

# Diffusion Model in Hyperspectral Image Processing and Analysis: A Review

Abstract
Hyperspectral image processing and analysis has important application value in remote sensing, agriculture and environmental monitoring, but its high dimensionality, data redundancy and noise interference etc. bring great challenges to the analysis. Traditional models have limitations in dealing with these complex data, and it is difficult to meet the increasing demand for analysis. In recent years, Diffusion Model, as an emerging generative model, has shown unique advantages in hyperspectral image processing. By simulating the diffusion process of data in time, the Diffusion Model can effectively process high-dimensional data, generate high-quality samples, and perform well in denoising and data enhancement. In this paper, we review the recent research advances in diffusion modeling for hyperspectral image processing and analysis, and discuss its applications in tasks such as high-dimensional data processing, noise removal, classification, and anomaly detection. The performance of diffusion-based models on image processing is compared and the challenges are summarized. It is shown that the diffusion model can significantly improve the accuracy and efficiency of hyperspectral image analysis, providing a new direction for future research.
Index Terms:
Hyperspectral image, diffusion model, High dimensional data processing, Noise removal, Data enhancement, Anomaly detection.
1School of Optical-Electrical and Computer Engineering, University of Shanghai for Science and Technology, No. 516, Jungong Road, Shanghai 200093, China (e-mail: huxing@usst.edu.cn, 242250440@st.usst.edu.cn, dwzhang@usst.edu.cn)
2School of Electronics and Electrical Engineering, Shanghai University of Engineering Science, Shanghai 201620, China (e-mail: dqq1019@163.com)
3Aerospace Information Research Institute, Chinese Academy of Sciences, Beijing 100094, China (e-mail: hongdf@aircas.ac.cn)
4School of Electronic, Electrical and Communication Engineering, University of Chinese Academy of Sciences, Beijing 100049, China.
I Introduction
Hyperspectral image processing and analysis have important application values in remote sensing, agriculture, environmental monitoring and other fields. With the advancement of technology, the resolution and spectral range of hyperspectral images are continuously improved, which provides the possibility of fine feature identification and classification. However, the high dimensionality, data redundancy, and noise interference etc. of hyperspectral images also bring great challenges to the analysis. Traditional processing methods have certain limitations in dealing with these complex data, so the introduction of new models and methods has become the focus of current research. In recent years, Diffusion Model (DM), as an emerging generative model, has shown unique advantages in hyperspectral image processing and analysis.
Hyperspectral imagery provides a wealth of spectral and spatial data by capturing spectral information of objects in multiple narrow bands. The high-dimensionality and complexity of this data make it widely used in agriculture, environmental monitoring, mineral exploration and other fields. However, hyperspectral image analysis also faces many challenges. First, there is a huge demand for high-dimensional data processing and storage with high computational complexity. Second, the high correlation between bands leads to data redundancy, which increases the difficulty of analysis. In addition, noise interference and mixed image element problems also affect the accuracy of the analysis.
Modeling using neural networks has been a dominant force over the last decade or so, and has had a huge impact in a variety of domains since the emergence of generative models, ranging from images [1, 2] and audio [3, 4] to text [5] and point clouds [6]. From a probabilistic modeling perspective, the key defining feature of a generative model is that it is trained in such a way so that its samples $\widetilde{{x}}\sim p_{\theta}(\widetilde{{x}})$ are from the same distribution as the training data distribution ${x}\sim p_{\theta}({x})$. Energy-based models (EBM) achieve this by defining non-normalized probability densities on the state space; however, these methods require Markov chain Monte Carlo (MCMC) sampling during both training and inference, which is a slow and iterative process [7] In the last few years, there has been a renewed interest in generative modeling due to the advancement of deep learning architectures in general, which have revealed improved visual fidelity and sampling speed. Specifically, Generative Adversarial Networks (GANs) [8], Variational Auto-Encoders (VAEs) [9] and Normalized Streams [10] among others, have emerged. In addition to this, generative models based on diffusion processes provide an alternative to existing VAEs, EBMs, GANs, and normalized streams, which do not require aligning a posteriori distributions, estimating tricky partition functions, introducing additional discriminator networks, or placing network constraints, respectively. To date, diffusion models have been found useful in a variety of domains, ranging from generative modeling tasks (e.g., image generation [11], image super-resolution [12] image restoration [13]) to discriminative tasks (e.g., image segmentation [14], classification [15], and anomaly detection [16]). Recently, the hyperspectral community has witnessed an exponential growth in the number of diffusion-based techniques, and a large number of studies have been dedicated to the application of diffusion models in various hyperspectral scenarios. Since diffusion modeling has recently received a great deal of attention from the research community, the literature has a large number of contributions in this direction. For this reason, this survey aims to provide a comprehensive overview of the latest advances and a comprehensive overview of such models in the field of hyperspectral spectroscopy.
Different from this paper [17], firstly, it is aimed at generalized remote sensing images, covering multimodal data such as multispectral, SAR, and panchromatic images, with a wider scope of research but lack of in-depth analysis of hyperspectral properties, and focuses on sorting out the technological paths of traditional tasks such as image generation and super-resolution reconstruction, with the contribution of systematically establishing a methodological mapping of diffusion modeling in remote sensing for the first time. Secondly, it focuses more on the panoramic depiction of diffusion modeling applications in remote sensing, and although the conditional diffusion model [11] as well as the latent diffusion model [18] are described, the exploration of the emerging architectures stops at early applications before 2023, lacks a systematic analysis of practical engineering constraints, and does not provide an in-depth discussion of the key technological challenges specific to hyperspectral imagery. Instead, this paper focuses on the specific modality of hyperspectral imagery, and designs a multi-perspective classification of diffusion models in hyperspectral domain from the core characteristics of high dimensionality of data and spectral redundancy etc., which provides a systematic taxonomy for the study of diffusion models and their applications. Meanwhile, this paper classifies the existing diffusion models into three categories: denoising diffusion probabilistic models (DDPM) [19] and score-based generative models (SGM) [20] and stochastic differential equation-based models (SDE) [21], and analyzes the use of diffusion models in hyperspectral domain with specific classification. In addition, we will systematically sort out the progress of the application of diffusion models in vertical tasks, which is more comprehensive than the direction of [17], and categorize the tasks into eight categories: data enhancement and synthesis, data generation and complementation, classification, segmentation, image generation, anomaly detection and noise suppression, and data recovery. The significance of this paper lies in the following: first, it fills the gap of systematic review of diffusion model in hyperspectral field; second, it elaborates the methodological innovation of adaptation in hyperspectral field more comprehensively; third, it reveals the unique challenges and optimization paths of diffusion model in hyperspectral scenarios; fourth, it realizes a leap forward breakthrough in the three dimensions of theoretical depth, technological breadth, and application value, and lays down a key theoretical and practical foundation for hyperspectral image processing to move from laboratory research it lays a key theoretical foundation for hyperspectral image processing to move from laboratory research to operational applications.
The overall structure of this paper is organized as follows: Section I will introduce the definition, characteristics, data properties, and main tasks in the analysis of hyperspectral images; Section II will introduce the theory, evolution, and variants of the Diffusion model, the analysis in comparison with the traditional model, and the application in image generation; Section III will introduce the use of Diffusion model in hyperspectral image analysis and its specific applications; Section IV will introduce the challenges and limitations of Diffusion model; Section V will introduce the comparison between Diffusion model and traditional methods, and conduct experiments and analysis; Section VI will provide a future outlook of Diffusion model in hyperspectral image; Section VII will summarize the content of this paper.
II Basics of hyperspectral imagery
II-A Definition and Characteristics of Hyperspectral Imagery
Hyperspectral image is a spectral image with spectral rate in the range of ${10}^{-2}\lambda$ order of magnitude, covering the range from visible to infrared light range of multiple continuous bands, through the hyperspectral sensors (imaging spectrometer) mounted on different platforms, it can be imaged in the ultraviolet, visible, near-infrared and mid-infrared regions of the electromagnetic spectrum by tens to hundreds of consecutively subdivided spectral bands in the target region, and each pixel point includes information of numerous spectral bands, forming a continuous spectral curve, which can obtain spatial information as well as spectral information at the same time, truly combining spectra and images.
Hyperspectral images benefit from the development and maturity of imaging spectroscopy technology, which is a sample that combines image information and spectral information, implying a wealth of features that are useful for feature identification and processing.
Image information is containing external quality information such as size, shape, defects, etc., and spectral information is containing different internal physical structure and chemical composition. Both the remote sensing image with multiple segments and the continuous spectral information of each image element can be obtained, realizing the multi-dimensional expression of feature characteristics, and the combination of spatial and spectral information, which is convenient for accurate identification and classification.
Hyperspectral images have a high spectral resolution, usually at the nanometer level, and are able to finely portray the reflectance spectra of features, differentiate the spectral characteristics of similar substances, and improve the ability to classify and identify features. At the same time, it has a large amount of information, redundancy of information, multi-segmented, data including dozens of hundreds of segments, huge amount of data, high requirements for storage and processing capacity, providing rich spectral information, able to capture subtle spectral differences, and correlate between them spatially and spectrally.
Finally, hyperspectral images are also characterized by fine spectral feature expression, nonlinear characteristics, low signal-to-noise ratio and high dimensionality of data. Moreover, hyperspectral images are widely used in many fields, such as agriculture, environmental monitoring, mineral exploration, medicine and so on, and can recognize and classify different substances.
Figure 1: Hyperspectral image (a) and Illustration of hyperspectral image (b)
II-B High-dimensional characterization of hyperspectral image data
The data of hyperspectral images are characterized by high dimensionality, which is mainly reflected in the multidimensional structure, data redundancy, computational complexity, and information richness.
Hyperspectral image is a three-dimensional data cube, containing spatial, radiometric and spectral dimensions, with a large number of bands, consecutive bands forming high-dimensional vectors, and covering a range of bands with narrow band widths and high resolution at the nanometer level, which finely reflects the physicochemical properties, while adjacent bands are highly similar, resulting in redundancy of information and increasing the complexity of the data. Hyperspectral image is a combination of spatial and spectral, forming a “map in one”, which makes the dimension of the information itself is higher, when analyzing the data, we have to consider the spatial up and down and spectral characteristics, increasing the complexity of the model. High-dimensional data requires a large amount of computational resources when processing, especially when performing tasks such as feature extraction, classification, segmentation, etc., the computational complexity increases significantly. The storage requirements are high due to the huge amount of data, which is in high demand and costly. Traditional machine learning methods are slow to train in high-dimensional space and are susceptible to noise interference, which is especially obvious in high dimensions and increases the difficulty of processing; although it can be automatically extracted under deep learning, a large number of markers are required and the cost is extremely high.
The high dimensional nature of the data creates excessive waste of resources for the processing, which can easily lead to dimensional disaster and inverted performance degradation. In this regard there is a great need for the use of dimensionality reduction methods, such as PCA, LDA, etc. to reduce the dimensionality, retain important information and reduce the complexity of the model. Under the complexity process, it is important to ensure that problems such as overfitting in training occur.
II-C Main tasks in hyperspectral image analysis
The main tasks of hyperspectral image analysis include classification, segmentation, and anomaly detection etc. These tasks have important application value in hyperspectral image processing and can help us to extract useful information from complex high-dimensional data. The following is a detailed description of these tasks:
Classification: Hyperspectral image classification refers to the assignment of each pixel or region in an image to a specific class. These categories can be different types of features, such as vegetation, water bodies, buildings, and so on. It is often applied to crop classification and pest detection in agriculture; land use classification and forest cover monitoring in environmental monitoring; and mineral type identification in mineral exploration. Often exists high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of classification; data labeling: the classification task usually requires a large amount of labeled data, and the labeling cost of hyperspectral images is high; mixed image element problem: due to the limitation of spatial resolution, one pixel may contain multiple features, which increases the difficulty of classification and other problems. The traditional methods are statistical-based methods (e.g., maximum likelihood classification), machine learning-based methods (e.g., support vector machine SVM). Deep learning methods are convolutional neural networks (CNN), recurrent neural networks (RNN), etc.
Segmentation: Hyperspectral image segmentation refers to dividing an image into multiple regions or objects, each with similar spectral characteristics. Segmentation can be pixel-level segmentation or object-level segmentation. It is often applied to the segmentation of tumor regions in medical image analysis, the segmentation of buildings and roads in urban planning, and the segmentation of crop regions in agriculture. Often exist high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of segmentation; noise interference: noise in hyperspectral images affects the accuracy of segmentation; mixed image element problem: a pixel may contain multiple features, which increases the difficulty of segmentation and other problems. The traditional methods are clustering based methods (e.g. Kmeans), edge detection based methods. Deep learning methods are Full Convolutional Networks (FCN), UNet, etc.
Target Detection: Hyperspectral image target detection refers to locating and identifying specific targets or objects in an image. These targets can be military targets, mineral resources, vegetation types and so on. Often used in military target identification and tracking; mineral exploration in the detection of specific minerals; environmental monitoring in the detection of specific vegetation or pollution sources. Often exist high-dimensional data processing: the high-dimensional characteristics of hyperspectral images increase the complexity of target detection; target diversity: the spectral characteristics of different targets may be similar, increasing the difficulty of detection; background interference: complex background will affect the accuracy of target detection and other problems. Traditional methods are based on spectral matching methods, statistical methods. Deep learning methods are Regional Convolutional Neural Networks (RCNN), YOLO (You Only Look Once) and so on.
Anomaly Detection: Hyperspectral image anomaly detection refers to the detection of pixels or regions in an image that are significantly different from their surroundings. These anomalies may be unknown targets or features. It is often applied to the detection of unknown targets in the military, the detection of pollution sources in environmental monitoring, and the early detection of pests and diseases in agriculture. Often exist high-dimensional data processing: the high-dimensional nature of hyperspectral images increases the complexity of anomaly detection; diversity of anomalies: different types of anomalies may have different spectral characteristics, increasing the difficulty of detection; background interference: complex background can affect the accuracy of anomaly detection and other problems. The traditional methods are statistical based methods (e.g. RX algorithm), clustering based methods. Deep learning methods are Autoencoder, Generative Adversarial Networks (GAN), and so on.
Noise suppression: Hyperspectral noise suppression refers to the reduction or elimination of noise in hyperspectral images by means of specific algorithms and techniques, in order to improve the quality of the image and the accuracy of the subsequent analysis. It is mainly used in agriculture, military, food industry, mining exploration and other fields. Commonly used methods are mainly frequency or transform domain frequency reduction, spatial frequency reduction, and deep learning frequency reduction, and so on.
Data recovery: Hyperspectral data recovery refers to the removal of noise and interference in hyperspectral images through specific algorithms and technical means to restore their original spectral and spatial information to improve the quality and usability of data, which belongs to the further implementation of noise suppression. The main application areas are agriculture, environmental monitoring, biomedicine, geological exploration and so on. The commonly used methods are spectral reconstruction, spectral dimension recovery, low-rank representation modeling, and deep learning methods, and so on.
Data Augmentation and Synthesis: Hyperspectral image data augmentation and synthesis refers to increasing the diversity and quantity of training data by generating new hyperspectral image samples, so as to improve the generalization ability of the model. It is often applied in model training to increase the diversity of training data and improve the generalization ability of the model; and in data completion to generate missing hyperspectral image data. Often exist high-dimensional data processing: the high-dimensional nature of hyperspectral images increases the complexity of data enhancement and synthesis; data quality: the generated samples need to be of high quality and diversity to ensure the validity of the model and other issues. The traditional methods are interpolation-based methods, statistics-based methods. Deep learning methods are Generative Adversarial Networks (GAN), Diffusion Model, and so on.
Data generation and completion: Hyperspectral data generation and completion is a core task in data science, aiming to solve the problems of missing data, noise interference and data scarcity, and at the same time extend the application scenarios through synthetic data. Data generation generates synthetic data that conforms to the real data distribution through algorithms, which is used to enhance the data scale or simulate specific scenarios; data completion fills in the missing values or repairs incomplete data to maintain the consistency and completeness of the data distribution. It is usually used in remote sensing, medical imaging, agricultural monitoring and other fields. The traditional methods are based on statistics (interpolation), machine learning (K-nearest neighbor, decision tree) deep learning methods such as GAN, VAE, Diffusion model, etc.
In summary, the main tasks of hyperspectral image analysis include classification, segmentation, target detection, anomaly detection and data enhancement and synthesis. These tasks have important applications in hyperspectral image processing and can help us extract useful information from complex high-dimensional data. Although these tasks face challenges such as high-dimensional data processing, noise interference, and mixed image element problems, the accuracy and efficiency of hyperspectral image analysis can be significantly improved by combining traditional and deep learning methods.
III Diffusion Model
Diffusion Model is an emerging generative model that has demonstrated powerful capabilities in the fields of image generation, denoising, and data enhancement in recent years. It gradually transforms noisy data into target data by modeling the diffusion process of data in time. The core idea of the diffusion model is to gradually transform the data into noise through the forward diffusion process, and then recover the original data from the noise through the reverse diffusion process. This approach performs well in high-dimensional data processing, image generation and denoising tasks, and especially has unique advantages in hyperspectral image processing.
III-A Theoretical Foundations of the Diffusion Model
Diffusion model is divided into two processes: forward and backward diffusion diffusion; and finally training objectives are performed to minimize the difference between forward and backward diffusion processes. The basic diffusion models are usually divided into (1) denoising diffusion probabilistic model (DDPM), (2) score-based generative model (SGM), and stochastic differential equation-based model (SDE/NCSN). The basic principles of these three diffusion models are described below.
III-A1 Denoising Diffusion Probabilistic Model (DDPM)
Denoising Diffusion Probabilistic Model (DDPM) [19, 22] utilize two Markov chains: a forward chain that perturbs the data into noise, and an inverse chain that converts the noise back into data. The former is usually manually designed to convert any data distribution into a simple prior distribution (e.g., a standard Gaussian distribution), while the latter Markov chain inverts the former by learning a transition kernel parameterized by a deep neural network.
New data points are generated by first randomly sampling a vector from the previous distribution and ancestrally sampling it via an inverse Markov chain [23].
Figure 2: Diffusion processes in diffusion models
Formally, given a data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$,, a forward Markov process generates a sequence of random variables $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ replaced by a transition kernel $q(x_{t}\mid{x_{t-1}})$. Using the probability chain rule and the Markov property, the joint distribution of $\mathbf{x}_{1},\mathbf{x}_{2}\ldots\mathbf{x}_{T}$ is later decomposed, conditional on $\mathbf{x}_{0}$, and denoted as $q(x_{1},\ldots{,x_{T}}\mid{x_{0}})$, which becomes:
$q(\mathbf{x}_{1},...,\mathbf{x}_{T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$ (1)
In DDPM, the transition kernel $q(x_{t}\mid{x_{t-1}})$ is drawn manually to incrementally transform the data distribution $q(\mathbf{x}_{0})$ into a processable prior distribution. A typical design of a transition kernel is a Gaussian perturbation, and the most common choice of transition kernel is:
$q(\mathrm{x}_{t}\mid\mathrm{x}_{t-1})=N(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\mathrm{I}).$ (2)
where $\beta_{t}\in(0,1)$ is the hyperparameter chosen prior to model training. This kernel is applied to simplify the discussion, although other types of kernels apply in the same vein. As observed by Sohl-Dickstein et al. 2015 [22] this Gaussian transition kernel allows us to marginalize the joint distribution in equation (1) to obtain an analytic form of $q(x_{t}\mid\mathbf{x}_{0})$ for all $t\in{0,1,\cdots,T}$. $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}{:}=\prod_{s=0}^{t}\alpha_{s}$, in particular:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right).$ (3)
Intuitively, this forward process slowly injects noise into the data until all structure is lost. To generate new data samples, the DDPM first generates an unstructured noise vector from the prior distribution (which is usually easy to obtain), and then gradually removes the noise from it by running a learnable Markov chain in the opposite time direction. Specifically, the inverse Markov chain is parameterized by the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ and the learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. We choose the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ because the forward process is constructed such that $q(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$. The learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ employs:
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\right).$ (4)
where $\theta$ denotes the model parameters and the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized by the deep neural network.
The key to the success of the network training goal is to train the reverse process $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$ to match the actual time reversal of the forward process $q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$. This is usually achieved by minimizing the KL scatter (Kullback-Leibler) between the two:
$\mathrm{KL}\left(q(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})|\||p_{\theta}(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})\right)$ (5)
=$\overset{(i)}{\operatorname*{=}}-\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})}[\log p_{\theta}\left(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T}\right)]+\mathrm{const}$ (6)
=$\begin{split}\overset{(ii)}{\operatorname*{=}}\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})}\left[-\mathrm{log}p(\mathbf{x}_{T})-\sum_{t=1}^{T}\log\frac{p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})}{q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})}\right]\\ +\mathrm{const}:=-L_{NLB}(\mathbf{x}_{0}).\end{split}$ (7)
$\overset{(iii)}{\operatorname*{\operatorname*{\geq}}}\mathbb{E}[-\log p_{\theta}\left(\mathbf{x}_{0}\right)]+\mathrm{const}$ (8)
where “const” is a constant that does not depend on the model parameters $\theta$. The goal of DDPM training is to maximize the variational lower bound or equivalently minimize the negative variational lower bound, and thus it can be efficiently estimated by Monte Carlo sampling [24] and efficiently optimized by stochastic optimization [25].
III-A2 Score-Based Generative Models (SGM/NCSN)
At the heart of Score-Based Generative Models (SGM/NCSN) [26, 20, 27] is the notion of a (Stein) score (aka fraction or score function) [28]. Given a probability density function p(x), the score function is defined as the gradient of the logarithmic probability density $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Unlike the Fisher score $\nabla_{\theta}\log p_{\theta}(\mathbf{x})$, which is commonly used in statistics, the Stein score considered here is a function of the data x and not of the model parameters $\theta$. It is a vector field pointing in the direction where the probability density function has the largest growth rate.
The key idea of Score-Based Generative Models (SGM) [20] is to perturb the data with a series of augmented Gaussian noises and jointly estimate the score function for all noisy data distributions by training a deep neural network model (called Noise Conditioned Scoring Network, NCSN, in [20]) conditional on the noise level. The samples are generated by linking the score functions with decreasing noise levels by using score-based sampling methods including Ronzmann Monte Carlo [29, 30, 31, 20, 21], stochastic differential equations [32, 21]], ordinary differential equations [33, 34, 35, 36, 21], and various combinations thereof [21]. In the formulation of the score-based generative model, training and sampling are completely decoupled, so that multiple sampling techniques can be used after estimating the score function.
[…]
$\mathbb{E}_{t\sim\mathcal{U}[[1,T]],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[\lambda(t)\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].$ (9)
$\begin{split}=\mathbb{E}_{t\sim\mathcal{U}\|1,T\|\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\sigma_{t}^{2}\|\nabla_{\mathbf{x}_{t}}\log q\left(\mathbf{x}_{t}\right)-\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].\end{split}$ (10)
=$\begin{split}\overset{(i)}{\operatorname*{=}}\mathbb{E}_{t\sim u[[1,T]]\mathbf{x}_{0}\sim q(\mathbf{x}_{0})\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}[\lambda(t)\sigma_{t}^{2}\\ \prod\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})-\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\parallel^{2}]+\mathrm{const}.\end{split}$ (11)
=$\begin{split}\overset{(ii)}{\operatorname*{=}}\mathbb{E}_{t\sim U[1,T],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\left\|-\frac{\mathbf{x}_{t}-\mathbf{x}_{0}}{\sigma_{t}}-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right\|^{2}\right]+\mathrm{const}.\end{split}$ (12)
$\begin{split}(iii)\\ =&\mathbb{E}_{t\sim\mathcal{U}[[1,T]],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I})}[\lambda(t)\|\epsilon+\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\|^{2}]\\ +\mathrm{const}.\end{split}$ (13)
where $\lambda(t)$ is a positive weighting function, $\mathbf{x_{t}}$ is computed from $\varepsilon$ and $\mathbf{x}_{0}$, $\mathcal{U}[[1,T]]$ is a uniform distribution on the set ${1,2,\cdots,T}$, and $\varepsilon$ is a deep neural network with parameter $\theta$ for predicting the given $\mathbf{x}_{t}$ and t noise vectors. Comparing the equations, the training objectives of the DDPM and SGM are equivalent once $\epsilon_{\theta}(\mathbf{x},t)=-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x},t)$ is set.
III-A3 Diffusion model based on stochastic differential equations (Score-SDE)
When DDPM and SGM can be further generalized to the case of infinite time step or noise level, where the perturbation and denoising process is the solution of a stochastic differential equation (SDE), a formulation known as Score SDE [21], which is utilized for the noise perturbation and sample generation, and the denoising process requires estimation of a score function of the distribution of noisy data.
The SDE was used to score the data using a diffusion process controlled by the following stochastic differential equation (SDE) [21]:
$\mathrm{dx}=\mathrm{f}(\mathrm{x},t)\mathrm{d}t+g(t)\mathrm{dw}.$ (14)
where f(x,t) and g(t) are the diffusion and drift functions of the SDE, and w is a standard Wiener process (aka Brownian motion) The forward processes in both the DDPM and the SGM are discretizations of this SDE. 2020 As demonstrated in the literature [21] by Song et al. the corresponding SDE for the DDPM is: $dx=-\frac{1}{2}\beta(t)xdt+\sqrt{\beta(t)}dw$. For the SGM the corresponding SDE is: $\mathrm{dx}=\sqrt{\frac{\mathrm{d}[\sigma(t)^{2}]}{\mathrm{d}t}}\mathrm{dw}$. where, $e\beta(\frac{t}{T})=T\beta_{t}$, $\sigma(\frac{t}{T})=\sigma_{t}$ when T tends to infinity.
Here $q_{t}({x})$ is used to denote the distribution of $\mathbf{x}_{t}$ in the forward process. For any diffusion process of the form of equation (14), Anderson proposed [42] to show that it can be inverted by solving the following reverse time SDE:
$\mathrm{d}\mathbf{x}=[\mathrm{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})]\mathrm{d}t+g(t)\mathrm{d}\overline{\mathbf{w}}.$ (15)
where $\bar{\mathbf{W}}$ is the standard Wiener process when time flows backwards and dt denotes an infinitesimal negative time step. The solution trajectories of this reverse SDE have the same marginal density as those of the forward SDE, except that they evolve along opposite time directions [21]. Intuitively, the solution of a reverse-time SDE is a diffusion process that gradually converts noise into data. In addition, the literature presented by Song et al. in 2020 [21] proved the existence of an ordinary differential equation (ODE), the probabilistic flow ODE, whose trajectory has the same margin as the reverse time SDE. The probabilistic flow ODE is given by the following equation:
$\mathrm{d}\mathbf{x}=\left[\mathrm{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})\right]\mathrm{d}t.$ (16)
Both the inverse time SDE and the probabilistic flow ODE allow sampling from the same data distribution because their trajectories have the same margins. If one knows the fractional function $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ for each time step t, one can derive its inverse-time equations, which can then be used to generate samples by numerical techniques such as annealed Ronzwan dynamics [20], numerical SDE solvers [32, 21], numerical ODE solvers [33, 34, 21, 36, 43]], and predictor-corrector methods (a combination of MCMC and numerical ODE/SDE solver combination) [21]. As with the SGM, the transient fraction model $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is parameterized and the fraction function is estimated by generalizing the fraction matching objective in Equation (13) to continuous time, resulting in the following objective:
$\begin{split}\mathbb{E}_{t\sim\mathcal{U}[0,T],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{0t}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)\|^{2}\right].\end{split}$ (17)
where $\mathcal{U}[0,T]$ denotes the uniform distribution on $\mathbf{[}0,T]$.
The above is an introduction to the principles and formulas of the basic diffusion model, through their theoretical support, which makes the diffusion model work in the modern image field, solving certain limitations of the traditional model, but also brings defects. The study of Denoising Diffusion Probabilistic Models (DDPM) is the main focus in targeting the hyperspectral field, and it is also the most applied diffusion model in this field, the other two types of applications are relatively less, and they will also be introduced in the following sections.
III-B Evolution of the Diffusion model with different variants
Diffusion model started in 2015, Sohl-Dickstein et al. first proposed the theoretical framework of diffusion model, which combines nonequilibrium thermodynamics with generative model for the first time [22].In 2019, Song Yang et al. proposed Score-Based Generative Models (SGM/NCSN) [20], combining Score-Based Generative Models with Score Matching,which is a score-based generative model that generates by estimating the logarithmic gradient of the data distribution. Immediately following the optimization of the model in 2020, the team proposed NCSN++ (Noise-Conditioned Score Networks++) [27], which improves the network architecture (introducing residual connectivity and adaptive normalization), optimizes the noise scheduling strategy, and proposes a more stable training objective function. In the same year, Ho et al. proposed DDPM (Denoising Diffusion Probabilistic Models) [19], which used fixed noise scheduling and reparameterization techniques and became the basis of modern diffusion models. Later, in order to address the deficiencies of DDPM in terms of generation quality, training efficiency, and log-likelihood optimization, and so on, Improve-DDPM (IDDPM) [44] was proposed by Nichol et al. in 2021, which introduced learnable variance, cosine noise scheduling, and hybrid loss function, and accelerated the sampling strategy by time-step-importance sampling, which made the DDPM effectively improved , but the computation is complicated, the dataset has limitations, the generalization ability is insufficient, and the generation quality and speed belong to this and that.Song et al. proposed DDIM (Denoising Diffusion Implicit Models) [35] in the same year, which accelerates the sampling through non-Markov chains. Meanwhile, Rombach et al. proposed LDM (Latent Diffusion Model) [18], which shifts the diffusion process to latent space instead of operating directly in pixel space, dramatically reducing the computational cost and ensuring the generation quality. The enhancement for NCSN++ was also introduced in the same year with Score-SDE (based on stochastic differential equations) [21] as well as ODE (probabilistic streaming) [21], which optimize the sampling process with stochastic and constant differentiation and optimize the sampling process, respectively. Also in 2021, Dhariwal and Nichol proposed Classifier Guidance (Conditional Diffusion Model/CDM) [11], is to significantly improve the quality of generation by incorporating a classifier to guide the generation.In 2022, Rombach et al. published Stable Diffusion [18], which is based on the LDM model improved to enhance text-to-image generation. There are important variants [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57] that have been improved since then, but do not constitute a “landmark” model. Below is a timeline of the evolution of the diffusion model and its main variants:
Figure 3: Evolution timeline of the Diffusion model
III-C Traditional Deep Learning Models
Generative Models (GM) is a class of machine learning models that can learn distributions from data and generate new samples. It includes data generation, data enhancement, image restoration, anomaly detection and so on. It is mainly valuable in computer vision, natural language processing and other fields. In this paper, we will focus on the traditional generative models: the GAN, VAE, autoregressive models.
Generative adversarial network (GAN) is another important network framework for unsupervised anomaly detection techniques.GAN was originally conceptualized in the zero-sum game framework of game theory. It was proposed by Ian Goodfellow et al. in 2014 [8]. The core idea is to generate data by adversarial training of two neural networks: a generator and a discriminator. The generator is responsible for generating data, and the discriminator is responsible for distinguishing the generated data from the real data. Through this adversarial process, the generator gradually learns to generate data similar to the distribution of the real data, which is fast to generate, but the training is unstable and prone to pattern collapse problem [58, 59, 60].
Variational Auto-Encoder (VAE) [9] is a generative model that combines the advantages of Autoencoder and probabilistic graphical models to learn latent representations from data and generate new data samples.The core idea of VAE is to approximate the latent distributions of the data by Variational Inference (VII), and to generate the decoder by generating new data with lower generation quality but stable training [61, 62, 63].
Autoregressive Models (ARMs) [64] are a class of generative models that model sequence data by assuming that the current data point is dependent on previous data points. The core idea of autoregressive models is to utilize the temporal dependence of sequence data to generate new data points step by step, e.g., RNN (Recurrent Neural Network) forms a recursive body in the process of processing the data; LSTM (Long Short-Term Memory Neural Network)) a kind of RNN with a special pattern. Autoregressive models generate the complete data by generating each part of the data step by step, which is slow in generating speed, but generating high quality [65, 66].
III-D Successful Application of Diffusion Modeling in Image Generation
Diffusion model has a wide range of applications, for images on the application began in 2020, mainly based on Denoising Diffusion Probabilistic Models (DDPM), first of all, the initial proposed DDPM [19], IDDM [44], and DDIM [35], etc., through the Markov chain as well as the addition of the noise and denoising process, and finally the prediction of the noise residuals to optimize the quality of the generation. However, the number of sampling steps is numerous and the computation speed is slow. Then ADM-G (Classifier-Guided-Diffusion) [11], introduced by OpenAI team in 2021, made the model surpass GAN in FID metrics for the first time by introducing classifier-guided generation. but dependent on classifiers and high computational cost. Later, in order to optimize text-image alignment and super-resolution generation, Imagen [46] was introduced by the Google Research team, which combines the T5-XXL text encoder with the cascade diffusion architecture, and has outstanding text alignment capability, but with great resource requirements. Later, in order to realize multimodality and efficient generation, Stabel-Diffusion [18] was proposed by Robin Rombach et al. in 2023, which compressed the diffusion process into the latent space, significantly reducing the computational complexity, but losing information in the latent space. Meanwhile, also by the OpenAI team, CLIP-Guded-Diffusion [49] was introduced, which combines the CLIP model to realize cross-modal generation, supports complex semantic control, and generates semantic consistency, but has large training requirements and slow generation speed. Subsequently, in order to unify the framework and extend the multi-tasking domain, OmniGen [53], the first unified framework supporting tasks such as text generation, image editing, and so on, without additional plug-ins, was introduced by the Wisdom Source Institute in 2024. Then, the Byte Jump team introduced D-Dit [55], which combines continuous and discrete diffusion techniques to achieve bidirectional text-image generation, support multimodal tasks, and improve parameter efficiency by about 20%. In 2025 by Yifan Zhou et al. proposed AF-LDM [57], which introduces shift isovariance loss, improves the UNet attention module, reduces potential spatial noise bias, improves video generation consistency, and enhances the model robustness, etc.
At the same time, the diffusion model will continue to be innovated and developed in the hyperspectral field in the future, and these models are improved on the basis of the main diffusion model, aiming at solving the problem of image in classification, anomaly detection and recovery, and so on.
Figure 4: Improved division of the diffusion model according to different variants
TABLE I: Basic Diffusion Models and Variants
Basic Diffusion Model Major Variants Reasons for Improvement Improvement Point Drawbacks Applications
DDPM IDDPM Improve generation quality and training stability. Introduction of cosine noise scheduling and hybrid objective function. to optimize the training process. Generation is still slow. Image generation, audio synthesis.
DDIM Accelerating the generation process. Non-Markov chain inverse process allowing large step size sampling. Sacrificing some generative diversity for speed. Fast image generation, interpolation.
Classifier-Guided Enhanced generation control. Using classifier gradients to guide the direction of generation and improve the quality of conditional generation. Additional training of classifiers is required, increasing computational cost. Conditional image generation, text-to-image.
LDM/Stable Reduced computational costs. Diffusion in latent space to reduce computational overhead in high-resolution images. Latent space compression may lose detail. High-resolution image generation, art creation.
SGM/NSCN NSCN++ Enhancement of model presentation skills. Introducing deeper network structure and improved score matching goals. High training complexity. Complex distribution modeling, scientific simulation.
Score-SDE Harmonization of diffusion frameworks. Modeling Continuous Diffusion Processes by Stochastic Differential Equation (SDE) for More General Theory. The sampling process is computationally intensive. Multimodal generation, cross-domain generation.
ODE Improved generation process stability. Replacing SDE with ordinary differential equations (ODE) for deterministic sampling. No significant increase in generation speed compared to SDE. Controlled generation, inverse problem solving.
IV Applications in Hyperspectral Image Processing and Analysis
Deep learning models are widely used in various fields and show powerful advantages, in recent years, some scholars have also applied them to the field of hyperspectral, while making many achievements. The application in hyperspectral image analysis is embodied in many aspects, mainly including data enhancement and synthesis, generation and complementation of hyperspectral data, image classification, segmentation, target detection, and anomaly detection, and many other aspects, in this section, the hyperspectral image processing is divided into two aspects: one is the image processing under the traditional deep learning model, and the other is the image processing under the diffusion-based model. The following is a description of these methods.
Figure 5: Deep learning model-based hyperspectral image processing
Figure 6: Timeline of the birth of each model
IV-A The main Diffusion model used in the field of hyperspectral imagery
Diffusion Model based Hyperspectral Image Processing, which was first applied in the field of Hyperspectral in 2022, meanwhile, it was obtained through the research that the main use of Diffusion Modeling methods in this field is mainly Denoising Diffusion Probabilistic Models (DDPM), and this paper also presents the technical analysis of DDPM and its variants in the field in this section. The main diffusion models for hyperspectral image processing based on diffusion modeling are DDPM [19], IDDPM [44], DDIM [35], LDM [18], Classifier-Gided [11], Stable-Diffusion [18].
First of all, the basic principles of DDPM have been introduced above, and the basic principles and formulas of IDDPM and DDIM are the same as those of DDPM, so we won’t repeat them here.LDM and Stable-Diffusion belong to the same category, in which the core ideas are the same, and it can be said that Stable-Diffusion is the specific implementation version of LDM, so the following will introduce the following LDM (Latent Diffusion Model), Classifier-Guided (Conditional Diffusion Model) two basic principles, formula.
IV-A1 Conditional Diffusion Model (CDM)
Conditional diffusion models are similar to the development of GANs in that diffusion models begin with unconditional generation, followed immediately by conditional generation [67]. Unconditional generation is often used to explore the upper limits of the model’s capabilities, whereas conditional generation is more beneficial for applications as it allows the output to be controlled according to the human’s wishes.
The first work to introduce this condition in diffusion models was [11], which guided the generation of diffusion models by adding classifiers to trained diffusion models, hence the term guided diffusion models. Although this approach has a low training cost, it increases the inference cost by using the classification results to guide the sampling process of the diffusion model and has poor control over the details to produce images with better results. Therefore, the Google team [68] decided to adopt a simple idea of controlling the generated results by conditionally retraining the DDPM and named it Classifier-Free Guidance.
Formally, given the conditional information $\mathbf{c}$, the distribution of the combined DDPM changes from equations (2)(3), and accordingly, the optimization objective and sampling process are modified to the following form:
$\mathcal{L}_{\mathrm{con}}(\theta)=\mathbb{E}_{x_{0},\epsilon,c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,c,t\right)\right\|^{2}\right].$ (18)
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\tilde{\alpha}_{t}}}\epsilon_{\theta}(y_{t},c,t)\right)+\hat{\beta}_{t}z.$ (19)
This approach is more widely used than guided diffusion models and is the basis for numerous models of important diffusion (e.g., DALL-E2 [45], Imagen [49], and stabilizing diffusion (SD) [18]).
IV-A2 Latent Diffusion Model (LDM)
The computational cost of direct manipulation of the diffusion model in the original image pixel space is high, Rombach et al [60] proposed to train and reason about the diffusion model in a low-dimensional latent space and named it Latent Diffusion Model (LDM).
Figure 7: Architecture of LDM, where E and D denote the encoder and decoder, respectively
As shown in Fig. 7, the LDM consists of three parts. First, the specific diffusion process of the LDM, which shifts the workspace from pixel space to latent space using a pre-trained autoencoder, is shown in the leftmost gray box in Fig. 7, where the noise is added to feature $\mathbf{z}$ to get the lower half of $\mathbf{z_{T}}$ as forward diffusion; the middle orange box is in latent space, where the U-Net with crosscutting attention supports the multimodal inputs in order to restore $\mathbf{z_{T}}$ to $\mathbf{z}$ as reverse diffusion. The right side box is the conditional encoder, which encodes various conditions (e.g., text and image) into feature vectors $\tau_{\theta}$ and feeds them into U-Net to generate an image.LDM is able to effectively reduce computational cost without degrading the quality of image synthesis while speeding up the inference speed, which is widely used in natural scenes and provides inspiration for the application of diffusion models to hyperspectral applications.
IV-B Application of Diffusion model in image processing
IV-B1 Data Enhancement and Synthesis
In data enhancement and synthesis, usually super-resolution techniques have an important role and unique advantages in improving quality. Super-resolution aims at reconstructing high-resolution images from low-resolution images [69], but the images can be disturbed, due to imaging equipment, weather conditions or downsampling. The loss of detail in hyperspectral images can be significant when compared to natural images, making inverted reconstruction of high-resolution images more challenging. Before the advent of diffusion models, deep neural network based approaches, e.g., Convolutional Neural Networks (CNNs) and Transformers, were finding suitable mapping functions to integrate feature information into a single step. To achieve this goal, methods including self-supervised learning [70], adaptive marker selection [71] and frequency information assistance [72] were used. However, the diffusion model-based approach fuses low-resolution images and bootstrap images in each step of the diffusion process, which helps to better integrate different image information and, at the same time, has gained wide attention in the hyperspectral field.
Although HSI has a high spectral resolution, its spatial resolution is relatively low [73], which may limit the performance of various HSI-based applications. To obtain high-resolution hyperspectral maps, there are two types of methods: panchromatic sharpening [74] and multispectral and HSI fusion [75]. The former enhances HSI by injecting detail information in the panchromatic (PAN) image, while the latter utilizes multispectral images to help hyperspectral maps learn spatial details. For example, Shi et al [76] used multispectral and hyperspectral images in tandem as conditional inputs to a diffusion model, and then generated hyperspectral maps with high resolution from the information found in the two image modalities.Qu et al [77] devised a conditional recurrent diffusion framework that allows the spatial SR process and the spectral SR process to be guided by the multispectral image and the hyperspectral map, respectively, to generate high-resolution hyperspectral maps with complementary spatial and spectral information to generate high-resolution hyperspectral maps.
More contemporary research focuses on the use of PAN image HSI super-resolution [78, 79, 80, 81, 82, 83, 84].Meng et al [79] proposed a modal calibration module to enhance and extract features in two images, where the enhanced features are used as conditional inputs to the diffusion model.Cao et al [78] argued that the different image modal unique information should not be mixed for processing and designed two conditional modulation modules to extract coarse-grained style information and fine-grained frequency information as conditional inputs.
Figure 8: Converged network fusion mode [82]
In order to take full advantage of the unique information of different modalities, Li et al [82] proposed a panchromatic sharpening network based on a biconditional diffusion model, where HSI and PAN are used as independent conditional inputs to learn the spectral features and spatial texture, respectively. The network structure is designed based on LDM and achieves excellent SR performance while significantly reducing the computational cost.Rui et al [84] also perform the sampling process in low-dimensional space. Based on the fact that the high-resolution hyperspectral map can be decomposed into the product of two low-rank tensors, one of the low-rank tensors is first computed using the low-resolution hyperspectral map. This tensor, together with the LRHSI and PAN, is then conditioned and fed into a pre-trained diffusion model [85] for the other low-rank tensor. Unlike the above methods, this method is a completely unsupervised deep learning method that does not require the involvement of high-resolution hyperspectral maps in any of the processes, providing feasibility for its application in practice.
IV-B2 Data Generation and Completion
Generation and Complementation in general can be categorized as text-to-image and image-to-image, but in the hyperspectral domain only image-to-image exists. For diffusion models, diffusion models have image bootstrapping to generate new images. Guided images can be in various forms such as maps [86], semantic layouts [87, 88, 89, 90]. Even though these images contain only some specific information, they still show excellent performance in image generation. For example, Espinosa and Crowley [86] proposed to use maps and historical maps to train conditional diffusion models, which generated more realistic satellite images.Yuan et al. [87] generated high-quality hyperspectral images under the guidance of semantic masks, and also solved the intrinsic problem that diffusion models require a long time of training for model convergence. However, these things alone are not enough, so it is important to generate not only high-quality images, but also image-annotation pairs that are more practical. For example, Zhao et al [89] implemented a two-stage fine-tuning of the SD model to enable the generation from noise to annotations and then from annotations to images.Toker et al [90] extended the standard diffusion model to a joint probabilistic model for images and their corresponding labels, which enabled the simultaneous generation of labels and images.
Hyperspectral images (HSI) can provide richer spectral spatial information. Meanwhile, the acquisition cost of HSI is much higher than that of multispectral images. Therefore, subsequent scientists then used easily available multispectral remote sensing images to generate hyperspectral images [91, 92]. However, the hyperspectral maps generated by the diffusion model used need to match the input noise dimensions with the spectral bands, which leads to too much noise sampling space, thus making the model convergence difficult. To solve this problem, Zhang et al [91] proposed a spectral folding technique to convert the input hyperspectral map into a pseudo-color image before training the diffusion model.Inspired by the LDM, Liu et al [92] used Conditional Vector Quantization GAN (VQGAN) [93] to obtain the Latent Diffusion Model latententent space for the hyperspectral map, and performed the diffusion model training and sampling in this space process. Afterwards, in order to address the status quo of high cost and low spatial resolution, a DM-based hyperspectral spectral decomposition method, DiffUn [94], was proposed by Deng et al. The spectral prior distributions learned by unconditional DM and the likelihoods derived from the LMM were combined to facilitate the deblurring performance and synthesize the hyperspectral maps.A novel ISPDiff [95] model was proposed by Dong et al. A transparent unified scale-flexible model was designed to solve such problems using an iterative optimization algorithm with MSRN for scale-matching recovery and CMUN for cross-scale upsampling.
Overall, image-to-image generation methods based on conditional diffusion models [67, 68], which use bootstrap images as conditional inputs to conditional DDPM, are essentially generating images from noise rather than directly. In order to realize a true image-to-image conversion, Yu et al. proposed UnmixDiff [96] to address the feasibility of depth generation model in HSI synthesis task, which synthesizes HSI from the perspective of material distribution and introduces a clear physical interpretation into the hyperspectral image synthesis task. By introducing an unmixing autoencoder in the diffusion generation model, the hyperspectral image synthesis is projected from the high-dimensional spectral domain to the low-dimensional abundance domain, which greatly reduces the computational complexity and effectively improves the spectral confidence of the generated hyperspectral images. By constructing a diffusion generation model in the abundance space, at the same time, hsi with real spatial details can be generated to maintain spectral consistency. Also at the same time this type of approach, which first employs self-supervised learning to extract the feature embeddings of the input patches, and then uses these embeddings as conditional inputs to guide the learning of the diffusion model. Notably, this preserves the spatial arrangement of each color patch in the original image so that the generated patches can be assembled into a large and coherent image.
Figure 9: UnmixDiff generates a simple diagram of the HIS [96]
IV-B3 Classification
Among the applications in the hyperspectral field, classification task is very important in image analysis as it can accurately identify and characterize different structures and anomalies in an image. It helps professionals in their ability to interpret large amounts of complex data Despite this potential, image classification aims at obtaining more valuable information about detailed features, such as buildings or grass, by assigning each pixel to a specific class [97, 98]. However, hyperspectral images contain many diverse and complex scenes and information, which increases the difficulty of accurate classification. Therefore, the current use of diffusion model to enhance classification is still a major challenge that needs to be further strengthened and solved. Meanwhile, because diffusion models can learn and simulate more complex data distributions than other deep learning models, more and more people are committed to applying diffusion models to hyperspectral image classification, which leads to better and better classification results [99, 100, 101, 102, 103, 104].
For image classification, hyperspectral images exhibit more complex data distributions, posing greater challenges to the application of traditional deep learning models in classification. The good thing, however, is that diffusion models can better capture the spectral-spatial joint features of hyperspectral images [105, 106, 107], improve the classification accuracy [108, 109], and improve the generalization ability [110]. Based on the diffusion model to realize the above properties, Zhou et al [111] used the temporal information of the diffusion model to build a time-stepping feature library, and also proposed a dynamic fusion module to combine spectral-spatial features with temporal features, and these improvements can make it possible to obtain sufficient image information before classification.Li et al [112] designed a two-branched diffusion model, which extracts features from HSI and LiDAR images, which achieves complementarity of information between different modal images, enhances the distinguishability between pixels, and also shows excellent performance in more complex hyperspectral tasks [37].Qu et al [113] go deeper by setting the encoder of the two-branch diffusion model to operate in parameter-sharing mode, which ensures that the shared features are extracted from the multimodal images. In addition, Chen et al [114] utilized the diffusion model to assist deep subspace construction and achieved excellent HSI classification performance in an unsupervised manner. Different from the above approaches, Qu et al [115] used a diffusion model to combine the SR (super-resolution) network and classification network to progressively reconstruct high-quality images by iteratively iterating the image super-resolution and classification tasks to achieve mutual guidance of the two tasks. The multi-scale classification results are utilized to guide the target-scale image to learn new knowledge and further reconstruct high-quality images. The super-resolution performance and classification performance are significantly improved.Jiang et al. proposed an image fusion [107] classification network based on Denoising Diffusion Probabilistic Models (DDPM), and after fusion, the classification accuracy is significantly improved.Zhu et al. proposed a Spatial-Spectral Diffusion Comparison Representation Network (DiffCRN) based on Denoising Diffusion Probabilistic Models (DDPM) [102], which is structured spatially self-attentive denoising module (SSAD) and spectral group self-attentive denoising module (SGSAD), and improves the efficiency of spectral spatial feature learning by designing new DDPM models loss and CL with log absolute error (LAE), as well as designing an adaptive weighted addition module (AWAM) and cross-time step spectral spatial fusion module (CTSSFM), which can effectively mitigate the interference due to spatial spectral heterogeneity and noise effects, etc., to effectively mitigate the interference due to spatial spectral heterogeneity and better classification.
Figure 10: Diffusion model based image classification process diagram
IV-B4 Segmentation
Image segmentation is an important task in computer vision, like semantic segmentation [116], image segmentation [14, 117], and in the field of deep learning, semantic segmentation is a subtask of image segmentation [112, 118]. Hyperspectral image segmentation, which investigates simplifying the complexity of an image by decomposing it into multiple meaningful image segments. Specifically, it facilitates image analysis by providing useful information about hyperspectral related fields. However, deep learning models typically require a large number of different pixel-annotated training data to produce generalizable results. Nonetheless, the number of images and labels that can be used for hyperspectral image segmentation is limited due to the time, cost, and expertise required. For this reason, diffusion modeling has emerged as a promising approach in image segmentation research that synthesizes labeled data and eliminates the need for pixel annotation data [102, 104].
Dong et al [95] are using Denoising Diffusion Probabilistic Models combined with Scale Matching Recovery Networks (P2MSRN) and cross-scale model-driven up-sampling networks.Chen et al [119] design image conditional mixing guides and combine them with Denoising Diffusion Probabilistic Models.Zhang et al [110] utilize Diffusion Probabilistic Models for labeling information.Zhu et al [120] utilize Variational Auto-Encoder as well as semi-supervised diffusion model for labeling annotation.Li et al. proposed a multidomain diffusion-driven feature learning network (MDFL) [112] to improve the pixel-level classification (i.e., segmentation) performance of high-dimensional remote sensing images through a multidomain diffusion model and feature reuse mechanism.Zhu et al. proposed a diffusion-based diffusion model coupled with contrast learning for diffusion contrast representation networks (DiffCRNs) [102] to extract the spatial-spectral features and finally generates pixel-level classification circles. Xu et al. proposed a data-enhanced semantic segmentation framework based on the diffusion model [104], which generates pseudo-labeled data through the diffusion model, and uses ControlNet to achieve cross-modal generation of “label→multispectral image”, and multi-stage training, which improves the segmentation accuracy.
IV-B5 Target Detection
Target detection differs from pixel-level ground classification in that target detection for hyperspectral images uses bounding boxes to localize instances like an airplane, a vehicle, or a ship in the image [121, 122]. One of the most challenging problems in target detection is that the training data is ambiguous, disturbed by external conditions [123, 118], and lacks realism and diversity. This problem is due to the intrinsic nature of hyperspectral images, which usually results in objects that need to be detected being complex, difficult to separate, and distributed in different regions of the image. Therefore, augmenting the objects to be detected with diffusion models has become an effective solution.
In order to solve all the above mentioned problems, Yu et al. proposed Diff-Mosaic (Enhanced Target Detection via Diffusion Prior), a diffusion model based approach [118], which introduces an enhancement network called Pixel-Prior, which coordinates pixels to generate highly coordinated and lifelike mosaic images, using a powerful diffusion model prior that resamples the pixels with samples to introduce real-world information. This resampling makes the generated samples realistic and diverse. Similarly, Qi et al. proposed a novel joint framework for underwater hyperspectral image restoration and target detection based on conditional diffusion framework [123], which uses a variable spectral group extraction module and a joint underwater hyperspectral image restoration and target detection (JURTD) module to simulate different ranges of underwater target spectra, achieve high-quality restored images and superior detection performance, and effectively overcome the challenges of underwater hyperspectral imaging. spectral imaging challenges.Zhu et al. proposed ODGEN (Domain-specific Object Detection Data Generation with Diffusion Models), a novel approach [124] to generate high-quality images conditioned on bounding boxes, thereby facilitating data synthesis for object detection.The use of synthetic visual cues with spatial constraints and textual descriptions by object to control the diffusion model allows ODGEN to perform well at demonstrate robustness when dealing with complex scenes and specific domains. Meanwhile, in order to improve the precision and accuracy, Wang et al. proposed OrientedDiffDet (pairwise object-oriented diffusion model) [125], which combines oriented frame representation, feature alignment and accelerated diffusion, a process that detects objects directly from a set of random frames. In the training phase, the horizontal detection boxes are first converted to oriented detection boxes. The model then learns to reverse this conversion process by diffusing from ground truth oriented boxes to a random distribution. In the inference phase, the model progressively refines a set of randomly generated boxes to generate the final output.
Figure 11: OrientedDiffDet target detection diagram with targets boxed in different colors in the detection diagram [125]
IV-B6 Anomaly Detection
Image anomaly detection is an important topic in computer vision, aiming to highlight anomalous regions of an image [115, 126, 127, 128]. In recent years, generative modeling has greatly influenced the queries for anomaly detection and shown promising results. Therefore, it will be explored in this section how the diffusion model, as the main generative model in anomaly detection, can be applied in hyperspectral collar anomaly detection. In fact, there are relatively few studies on hyperspectral image anomaly detection based on diffusion modeling, however, there are relatively numerous applications in fields such as medical impact areas [129, 130], industrial images [131, 132], etc., so with the development of diffusion modeling in this hyperspectral image anomaly detection will be increasingly used.
Meanwhile, in the field of anomaly detection, it is often faced that on the one hand, the complexity of the background of the input hyperspectral image leads to the confusion of anomalies in the background. On the other hand, the lack of labeled samples for hyperspectral images leads to poor generalization of existing anomaly detection methods. To solve this problem, there Ma et al. proposed a novel solution for anomaly detection, BSDM (Background Suppression Diffusion Model) [133], which simultaneously learns the Latent Background distribution and generalizes it to different datasets to suppress the complex background. It is characterized by three aspects:(1) Pseudo background noise is designed for the complex background of HSIs and Diffusion Model (DM) is used to learn the latent background distribution therein. (2) For the generalizability problem, a statistical bias module is applied to adapt the BSDM to datasets with different domains without labeling the samples. (3) To achieve background suppression, the inference process of DM was innovatively improved by feeding the original hyperspectral image into a denoising network, which removes the background as noise. They were also the first to apply diffusion modeling to hyperspectral image anomaly detection. Wu et al. proposed the diffusion background dictionary method (DBD) [134], which combines the advantages of model-driven and data-driven based on the characteristics of HSI and HAD tasks.DM is intrinsically combined with the low-rank representation (LRR) model, and DM is utilized to obtain the key background dictionary tensor in the tensor LRR. Thus, anomalies can be accurately detected. According to the idea of RX algorithm in anomaly detection, the multivariate normal distribution approximating the HSI background is diffused to make it more suitable for suppressing the background. Also hyperspectral image anomaly detection is often faced with the lack of a priori information, essentially the challenge is to accurately reconstruct the background region while inferring the latent background of the anomalous region, due to which Chen et al. proposed a novel approach called Double Window Spectral Diffusion (DWSDiff) [56], which modeled accurate background estimation through an iterative spectral diffusion and reverse diffusion process. A dual-window strategy is also introduced to mitigate the effect of anomalous extended regions in the neighborhood on the background estimation. An anomaly generation strategy based on Principal Component Analysis (PCA) and Linear Spectral Mixture Model (LSMM) is devised to create realistic training data that overcomes the scarcity of labeled hyperspectral data. The same effect is improved.
Figure 12: Various diagrams under BSDM from left to right: anomaly diagram, 1D t-SN diagram, 2D t-SNE scatter diagram [133]
In conclusion, the road of diffusion model in hyperspectral image anomaly detection is still long and the research prospect is very great, more and more researchers will put their eyes here to promote the vigorous development of the field of anomaly detection.
IV-B7 Noise Suppression
The main challenge in hyperspectral imagery is to acquire images without losing important information. During the acquisition and/or further processing stages, the acquired image may be corrupted by noise or artifacts. Noise degrades the image quality and is especially noticeable when the objects are small and the contrast is relatively low. Due to the nature of the generative model, the diffusion model facilitates the solution of various noise reduction problems [135, 136, 137]. Among them, [136] is denoising in image restoration, where the robustness of denoising is enhanced by designing a multi-shape spatial rectangular self-attention module, a spectral latent diffusion enhancement module, and also by utilizing the diffusion model, which allows for a tighter alignment with the desired HSI.
In image processing, hyperspectral images are always accompanied by various noises due to the inherent limitations of imaging techniques and environmental conditions. In essence, the learning process of diffusion models is equivalent to the denoising process [19], which makes it possible to achieve excellent denoising performance in hyperspectral contexts. Different image modalities face different noise challenges. For example, hyperspectral image (HSI) recovery, which may be quite susceptible to distributional biases, is data-poor [138]. HSI needs to take into account the inhomogeneous distribution of noise in the spectral dimension [139] as well as the correlation of noise between different frequency bands [140, 141]. Fortunately, these noises can be efficiently removed by a diffusion model.Ceng and Cao et al. proposed Diff-Unmix [142] to address the challenge of recovering noise-degraded HSI through the fusion of spectral unmixing and conditional abundance generation, which employs a learnable block-based spectral unmixing strategy supplemented by a purely transformer-based backbone. We then introduce a self-supervised generative diffusion network to augment the abundance maps from the spectrally unmixed blocks. This network reconstructs noise-free decomposition probability distributions, effectively mitigating noise-induced degradation within these components. Finally, the reconstructed HSI is reconstructed by de-mixing reconstruction by mixing the diffusion-adjusted abundance maps with the spectral end-elements. he et al [140] proposed a truncated diffusion model that starts denoising from the intermediate step of the diffusion process instead of the pure noise in order to avoid destroying the inherently valid information in the HSI. In addition, Deng et al. proposed RGB image-enhanced hyperspectral Denoising Diffusion Probabilistic Models [143], which improved DM-based HSI denoising by combining RGB-DM pre-training on additional RGB images, and a fusion operator was introduced to integrate HSI-DM and RGB-DM, which improves the performance of DM-based HSI denoising.Kai et al. proposed a multi-scale spatial spectral for hyperspectral image frequency joint denoising network (MSFDN) [137], in order to describe and extract the noise more accurately and obtain better denoising effect, they constructed a subspace projection network to convert the hyperspectral image into a low-dimensional subspace, designed a noise separation network, a subspace inverse projection network and utilized the conditional diffusion model for better robustness.
Figure 13: Process diagram of noise suppression based on diffusion model
IV-B8 Data Recovery
Hyperspectral image (HSI) recovery aims to recover clean images from degraded observations [144], and the recovery of hyperspectral images plays a pivotal role in subsequent hyperspectral image applications. Hyperspectral data recovery faces challenges in effectively exploring the spatial nonlocal self-similarity and spectral low-rank properties inherent in HSI, while hyperspectral exists data scarcity. Existing model-based approaches have limitations in accurately modeling complex image features with manual a priori, and deep learning-based approaches have poor generalization capabilities, complex models, poor results, and parameter redundancy, Diffusion models, due to their impressive image recovery performance [145, 146, 147, 148, 149, 150, 151]. In particular, it has recently attracted attention in terms of noise robustness [92, 95, 140, 152, 115, 119]. Diffusion methods, however, are trained on large amounts of training data and perform very well within distributions, but may be quite susceptible to distributional bias.
To address the above problem, Miao et al. proposed a self-supervised diffusion model for HSI recovery, the Denoising Diffusion Probabilistic Models for Spatial Spectroscopy (DDS2M) [138], which works by inferring the parameters of the proposed variational spatial spectral module (VS2M) during reverse diffusion, using only the degraded HSI without any additional training data. A loss function based on variational inference is customized to enable untrained spatial and spectral networks to learn a posteriori distributions that serve as transitions in the sampling chain to help reverse the diffusion process. Chen et al [56] use a dual window, iterative diffusion model, and reverse reconstruction to mitigate the impact of background estimation. Li et al. propose an untrained diffusion model with a pre-trained diffusion model (HIR-DIFF) as a A supervised HSI recovery framework [141], which can recover clean HIS from the product of two low-rank components, is proposed based on singular value decomposition (SVD) and rank-revealed QR (RRQR) decomposition, and a novel exponential noise schedule is proposed to speed up the recovery process. Li et al. propose a spectral latent diffusion enhancement module [136], which introduces a latent diffusion-enhanced rectangular transformer for HSI recovery. A multi-shape spatial rectangular self-attentive module is proposed, which leads to a tighter alignment with the desired HSI.
Meanwhile, there are also data recovery problems like noise mixing characterized by signal correlated and signal uncorrelated components, visual perception is affected and discriminative image distortion. Si et al. proposed a new model of CASSIDiff to reconstruct CASSI metrics [153], which employs a feature fusion mechanism based on the Discrete Wavelet Transform (DWT) and a spatial-spectral attention mechanism, and has a high degree of reliability in the recovery task. reliability. Yi et al. proposed an efficient diffusion model (EDiffSR) [149] for RSI SR, developed an efficient activation network (EANet), and developed a practical conditional prior enhancement module (CPEM), which reduces the computational budget and retains more information. Meng et al. proposed a conditional diffusion model-based SRSISR method (FastDiffSR) [148], which reduces the number of sampling steps, employs residual images, and combines channel attention and spatial attention to obtain better visualization with significantly improved generation quality.
Figure 14: Diffusion model-based research progress in hyperspectral image processing and analysis the tasks are established in nine subfields: (1) data enhancement and synthesis, (2) data generation and complementation, (3) classification, (4) segmentation, (5) anomaly detection, (6) target detection, (7) noise suppression, and (8) data recovery. No abbreviated name, with two initial letters at the beginning and end
V Challenges and Limitations of the Diffusion Model
V-A Computational complexity and Time cost of the model
Diffusion Models achieve data generation and denoising through multi-step iterative Markov chains, whose computational complexity increases exponentially with the data dimensions. Taking the denoising diffusion probabilistic model (DDPM) [19] as an example, hundreds of forward and reverse diffusion computations are required to generate a single hyperspectral image, especially when processing hyperspectral data above 200 bands, and the GPU memory occupancy can be more than 20 GB (e.g., 1024 × 1024 resolution scenes). Although the improved DDIM [35] reduces the number of sampling steps to less than 50 through non-Markov chains, the spatial-spectral three-dimensional nature of hyperspectral data (see Section II.B) still leads to a single inference time as long as several minutes, which makes it difficult to satisfy real-time demanding tasks, such as agricultural monitoring and disaster warning. Although the Latent Diffusion Model (LDM) [18] reduces the computational volume by compressing the latent space, the information compression in the spectral dimension may lose key features (e.g., mineral absorption peaks), which affects the classification and anomaly detection accuracy (see Table II).
V-B Impact of high dimensionality of hyperspectral data on model training
The performance of diffusion models is highly dependent on large-scale labeled datasets, and the cost of labeling hyperspectral data far exceeds that of traditional RGB images. For example, field sampling is required to validate crop classes in agricultural applications, and the cost of labeling a single image can be hundreds of dollars. Although DDPM [19] generates synthetic samples through data augmentation, its fidelity is limited by the distribution of training data, and it is prone to Mode Collapse in scenarios with fewer samples (e.g., rare mineral detection). Experiments show that when the training samples are less than 1000, the classification accuracy of the diffusion model decreases by about 12% (F1-score from 0.85 to 0.75) compared to Variational Auto-Encoder (VAE) [9]. In addition, conditional diffusion models based on classifier guidance (e.g., ADM-G [11]) need to pre-train high-precision classifiers, but the class imbalance problem of hyperspectral data (e.g., less than 1% of anomaly samples in anomaly detection) leads to difficulties in classifier training, which restricts the model generalization ability.
V-C Adaptability issues to large-scale data
Real hyperspectral images often contain non-Gaussian noise (e.g., sensor thermal noise, atmospheric scattering noise), and most of the existing diffusion models are based on additive Gaussian noise assumptions (see Section III.A.1) The forward process design of the DDPM [19] is only applicable to isotropic Gaussian noise, and the denoising effect is significantly degraded for streak or PoissonGaussian hybrid noise (PSNR reduction of 8 15 dB). Although Score-SDE [21] extends the scope of noise modeling through stochastic differential equations, its inverse process relies on the accurate estimation of the score function, which is susceptible to error accumulation in bands where the signal-to-noise ratio (SNR) fluctuates drastically (e.g., near-infrared bands). For example, the Diffusion Models’ denoising performance degrades by about 20% compared to the conventional wavelet approach under extreme conditions where the noise standard deviation exceeds 30 dB. In addition, the mixed image element problem (see Section II.C) leads to a high degree of spatial coupling between the noise and the target signal, further exacerbating the modeling difficulty.
V-D Limitations of current models in practical applications
The black-box nature of diffusion models exposes them to credibility challenges in hyperspectral critical applications. For example, in anomaly detection tasks, the model may misclassify rare features as noise by overfitting the training data (literature [16]). Experiments have shown that LDM [18] based anomaly detectors have a 18% higher false alarm rate than the traditional RX algorithm (see Section II.C) in the detection of unknown mineral types. In addition, although the Latent Diffusion Model improves the computational efficiency, the physical meaning of the latent vectors is unclear, which makes it difficult for users to trace the generation logic of the detection results. In scenarios that require strict validation, such as military target identification, the interpretability disadvantage of diffusion models significantly constrains their practical application. Comparative studies have shown that traditional methods (e.g., SVM) [154] can provide a decision basis through feature importance analysis, whereas diffusion models can only output probability estimates and lack a transparent inference chain.
In summary, the practical deployment of diffusion models in hyperspectral image processing is still limited by core bottlenecks such as computational efficiency, data dependency, noise robustness and interpretability. Future research needs to combine hyperspectral data characteristics (e.g., spectral continuity, spatial-spectral correlation), develop lightweight architectures and adaptive noise modeling methods, and explore semi-supervised learning frameworks to reduce annotation costs, so as to promote its large-scale application in remote sensing, agriculture and other fields.
VI Comparative and Experimental analysis
In order to effectively illustrate the superiority of the diffusion model in processing hyperspectral images, this paper takes the experimental results of classification, target detection, anomaly detection, image recovery, and noise suppression as an example, and evaluates the performance of the diffusion model and other existing techniques through the visualization results and quantitative indexes.
TABLE II: Comparison of Diffusion Model for Hyperspectral Images with Traditional Deep Learning Models
Deep Learning Model Advantages Disadvantages Future Development Trends
Diffusion Model Generates high-quality image details. Specializes in denoising and data reconstruction. Suitable for dealing with high-dimensional features of hyperspectral images. High computational cost. Slow training and inference. Sensitive to hyperparameters. Optimizing computational efficiency. Combining physical models to improve reconstruction accuracy. Lightweight design to adapt to real-time processing needs.
Generative Adversarial Network (GAN) Generate realistic hyperspectral data to support data augmentation. Enhance robustness by adversarial training in classification tasks. Training is unstable (pattern crashes). Generating data may have artifacts. Improved training stability (e.g., Wasserstein GAN). Combined with attention mechanisms to enhance local feature generation.
Variational Auto-Encoder (VAE) Effective degradation and feature extraction. Generating latent space representations. Supporting unsupervised learning. Generated images are fuzzy and do not retain enough detail. limited ability to model complex distributions. Improved decoder structure for enhanced detail generation. Combined with diffusion model to improve generation quality.
Autoregressive Models (AR) Strong sequence modeling capability for pixel-by-pixel prediction. Excellent performance in temporal hyperspectral analysis. High computational complexity (cannot be parallelized). Difficult to handle large size hyperspectral images. Introducing sparse attention mechanism to reduce computation. Combining with Transformer architecture to improve the efficiency of long sequence modeling.
VI-A Experimental results and Performance assessment and Various metrics
VI-A1 Introduction to evaluation metrics and Datasets
The receiver operating characteristic curve (ROC) is a qualitative analysis of the model. The horizontal coordinate of the curve is the false positive rate (FPR) and the vertical coordinate is the true positive rate (TPR).
AUC (area under ROC curve) is the area under ROC curve, which is a quantitative analysis of the model. In practice, AUC is often used as a metrics for evaluating the accuracy of a method. The larger the AUC value, the better the detection performance of the algorithm, with mAUC representing the average AUC. where larger ${\rm AUC}_{(D,F)}$and ${\rm AUC}_{(D,\tau)}$ indicate better detection performance, and are combined together defined as ${\rm AUC}_{TD}$; ${\rm AUC}_{BS}$ represents the ability of the detector to minimize false alarms due to background interference; ${\rm AUC}_{ODP}$ denotes the overall probability of detection; ${\rm AUC}_{TD-BS}$ specifically evaluates the performance of the detector in recognizing a targets in background noise, taking into account both detectability and background suppression; ${\rm AUC}_{SNPR}$ reflects the detector’s ability to distinguish between signal and noise.
OA denotes the ratio of the number of correctly classified samples to the total number of samples in all test samples; AA denotes the average of the classification accuracy for each category; k × 100 is the Kappa coefficient is a measure of the consistency of the classifier, which takes into account stochastic consistency and provides a measure of the consistency between the classifier and the ground truth.
IoU is a measure of the degree of overlap between the predicted frame and the real frame; $\mathbf{P}_{d}$ is defined as the ratio of the number of correctly detected targets to the total number of targets; and $\mathbf{F}_{a}$ is defined as the ratio of the number of false alarms to the total number of pixels in the image.
PSNR is the Peak Signal to Noise Ratio, a measure of the difference in quality between a super-resolution processed image and the original high-resolution image.FID (Fréchet Inception Distance) is a metrics used to assess the quality of the generated images, measuring their similarity to each other. Patch Similarity evaluates the perceptual difference between the generated image and the original high-resolution image. parmas is the number of parameters for model training. entropy is usually used to measure the uncertainty and complexity of image information. niqe is the Natural Image Quality Evaluator, which evaluates the image quality metrics. SSIM (Structural Similarity) measures the similarity between two images, comparing images in terms of brightness, contrast and structure; FSIM (Feature Similarity) evaluates the image quality by calculating the phase coherence and gradient similarity in localized regions of the image, and is an improved version of SSIM; SAM (Sum of Absolute Errors) is another measure of the effectiveness of image denoising, which calculates the sum of the absolute value of the pixel differences between a denoised image and the original image counterparts.
TABLE III: Summary of datasets used for the following experiments
Datasets Sensor Spectral Band Size Resolution
Salinas AVIRIS 204 512$\times$217 3.7 m
Pavia University ROSIS 103 610$\times$340 1.3 m
Indian Pines AVIRIS 200 145$\times$145 20 m
SIRST FLIR LWIR 512$\times$512 0.5 m
NUDT-SIRST Infrared Sensor LWIR 640$\times$512 0.3 m
Airport-Beach-Urban AVIRIS 205/188/191 150$\times$150 20 m
Potsdam ISPRS RGB(3) 5000$\times$5000 5 cm
Toronto Toronto RGB(3) 1920$\times$1080 0.2 m
UC Merced USGS RGB(3) 256$\times$256 0.3 m
CAVE Hyperspectral Camera 31 512$\times$512 N/A
CAVE-Toy Hyperspectral Camera 31 512$\times$512 N/A
KAIST Thermal Camera LWIR 640$\times$480 0.5 m
VI-A2 Experimental results
In this section, several typical methods from the above methods will be selected and compared to study the performance of detection, where traditional methods:GAN, VAE, autoregressive models, other deep learning methods, are compared with diffusion models. Since the experimental dataset of each algorithm is different, a common dataset is chosen for comparison as much as possible, in order to verify the anomaly detection performance of different algorithms on the same dataset in a more reasonable way. Where, the bolded indicates the algorithm’s performance optimal value for this dataset.
Comparison of classification experimental results, as shown in Figure 15, the classification effect of the first three methods is very poor, different classification color adulteration, the latter three methods overall classification effect is better, among them, the diffusion model-based method SpectralDiff classification effect is the best, followed by DMVL, their most boundary delineation is better versus the other methods, with less color adulteration, and SpectralDiff, despite the tiny classification there are some problems, but the overall effect is significantly better than other mainstream methods:
Figure 15: Comparison of different datasets on classification methods for hyperspectral images: (a) Pseudo-color image (b) Real labeled image (c) SF [155] (d) miniGCN [156] (e) SSFTT [157] (f) DMVL [158] (g) SSGRN [159] (h) SpectralDiff [108]
TABLE IV: Comparison of Metrics for Different Classification Methods on Indian Pines Dataset
Methods Evaluation metrics (Indian Pines)
OA AA K×$\times$100 Time(s)
SF 67.782 80.639 63.820 29.81
miniGCN 63.916 69.091 59.439 0.95
SSFTT 91.566 95.661 90.392 31.58
DMVL 90.487 81.311 89.216
SSGRN 92.336 96.099 91.254 1.13
SpectralDiff 93.146 96.438 92.175 48.53
Table IV shows that the Spectral-Diff method is in the first place in Indian Pines dataset classification in terms of overall accuracy, average accuracy, and kappa coefficient, which is significantly improved over other methods and is the current optimal method.OA and AA are higher than the second place (SSGRN) by 0.8% and 0.34%, which indicates that the method is more robust in classification and more effective in capture features and reduce noise interference. However, it can also be seen that Spectral-Diff consumes much more time than other methods, which is its shortcoming.
TABLE V: Comparison of Metrics for Different Classification Methods on Pavia University Dataset
Methods Evaluation metrics (Pavia University)
OA AA K×$\times$100 Time(s)
SF 81.511 84.942 75.795 163.34
miniGCN 82.355 84.037 76.620 2.40
SSFTT 93.667 93.448 91.617 110.08
DMVL 89.768 86.116 86.580
SSGRN 93.850 94.436 91.919 4.99
SpectralDiff 94.775 93.843 93.061 256.23
As can be seen in Table V, the method has two metrics on the Pavia University dataset that are optimal 0.98% and 1.23% higher than the second (SSGRN), and the other is only 0.63% lower than the optimal, which is a relatively small gap, and it is still the optimal method among these methods, but the shortcoming of the use of samples is the long time consumed, which is more than twice as long as the other methods.
TABLE VI: Comparison of Metrics for Different Classification Methods on Salinas Dataset
Methods Evaluation metrics (Salinas)
OA AA K×$\times$100 Time(s)
SF 88.248 93.263 86.973 141.30
miniGCN 88.181 94.279 86.823 2.13
SSFTT 95.789 98.272 95.322 55.96
DMVL 97.005 95.853 96.668
SSGRN 96.539 96.354 96.144 2.98
SpectralDiff 98.971 99.465 98.854 242.16
Table VI shows the classification on Salinas dataset and the same conclusion as the previous two tables, combined with Tables IV, V, and VI, the Spectral-Diff method is optimal in terms of performance compared to the other methods, but it is more time consuming and slower. SSGRN is optimal if balancing accuracy and efficiency, and miniGCN is optimal if low latency and high tolerance accuracy.
Comparison of target detection experimental results, as shown in Figure 16, the overall effect of these mainstream methods used in target detection is good, some methods exist undetected certain points or misdetection, among them, the diffusion model-based method Diff-Mosaic detection effect is the best, for the detection of the point of the edge of the border details of the best description, and the real labeling map is basically consistent, basically there is no missing parts:
Figure 16: Comparison of different datasets on target detection methods for hyperspectral images: (a) Original image (b) Real labeled image (c) ALC-Net [160] (d) ACM [161] (e) UIU-Net [162] (f) DNA-Net [163] (g) Diff-Mosaic [118]
TABLE VII: Comparison of Metrics for Different Target Detection Methods on SIRST Dataset
Methods Evaluation metrics (SIRST)
IoU $P_{d}$ $F_{a}$
ALC-Net 70.00 71.71 23.91
ACM 76.17 86.31 16.07
UIU-Net 72.03 98.10 26.15
DNA-Net 76.97 95.41 3.54
Diff-Mosaic 79.44 99.99 3.19
Table VII shows that the Diff-Mosaic method has the optimal intersection ratio, detection probability, and false alarm rate for target detection on the SIRST dataset, which is 3.11% higher than the second place (DNA-Net), 1.89% higher than the second place (UIU-Net), and 87.8% lower than the second place (UIU-Net), respectively. Among them, the detection tends to 100% detection is very good with less error and is the best among these methods.
TABLE VIII: Comparison of Metrics for Different Target Detection Methods on NUDT-SIRST Dataset
Methods Evaluation metrics (NUDT-SIRST)
IoU $P_{d}$ $F_{a}$
ALC-Net 72.13 78.81 21.22
ACM 71.11 85.13 22.51
UIU-Net 89.00 98.73 6.02
DNA-Net 88.38 97.99 4.04
Diff-Mosaic 91.18 99.47 1.91
Table VIII shows the Diff-Mosaic in manual synthesis dataset NUDT-SIRST utilization, which is consistent with the metrics effect shown in Table 7, and moreover, it shows the superiority of this method among these methods with the best performance.
TABLE IX: Comparison of Parameters and Reasoning for Different Target Detection Methods
Methods Calculate the number of parameters Time (s)
ALC-Net 0.38 40.93
ACM 0.29 18.53
UIU-Net 4.7 43.42
DNA-Net 50.54 33.98
Diff-Mosaic 4.7 43.42
Table IX shows that Diff-Mosai is less in the number of parameters, but not the best, while the inference time is the slowest among them, showing the shortcomings of the method in applications, 57.3% higher than the fastest time (ACM). Meanwhile, combining Tables VII, VIII and IX, again Diff-Mosaic is the best if pursuing the best performance, DNA-Net is the best if balancing performance and efficiency, and ACM is the best if high tolerance and low latency.
Comparison of the results of the anomaly detection experiments, as shown in Figure 17, compared with the traditional mainstream methods, the diffusion model-based methods BSDM, DWSDiff in the anomaly detection of the best detection effect, high detection rate, low false detection rate, and the real labeling map is basically consistent with the other mainstream detection methods to detect the detection of the detection effect is poor, the detection rate is low (the existence of the detection of the anomalies do not exist), the false detection rate is high:
Figure 17: Comparison of different datasets on hyperspectral image anomaly detection methods: (a) Pseudo-color map (b) Real label map (c) GAED [164] (d) RGAE [165] (e) Auto-AD [166] (f) DFAN-HAD [167] (g) DirectNet [168] (h) BSDM [133] (i) TGFA [169] (j)DWSDiff [56]
TABLE X: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{(p,r)}$ $mAUC_{(p,r)}$ $mAUC_{(p,r)}$
GAED 0.90756 0.15556 0.02072
RGAE 0.87615 0.14519 0.02401
Auto-Net 0.95028 0.15303 0.00627
DFAN-HAD 0.96552 0.45211 0.09603
DirectNet 0.95193 0.12799 0.00484
BSDM 0.94486 0.17342 0.03055
TGFA 0.92655 0.14559 0.00629
DWSDiff 0.97427 0.55329 0.01635
TABLE XI: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{(p,r)}$ $mAUC_{gs}$ $mAUC_{opp}$
GAED 1.06312 0.8868 1.04240
RGAE 1.02135 0.8521 0.99734
Auto-Net 1.10331 0.9440 1.09704
DFAN-HAD 1.41762 0.8694 1.32160
DirectNet 1.07993 0.9470 1.07508
BSDM 1.11828 0.9143 1.08773
TGFA 1.07214 0.9292 1.06585
DWSDiff 1.52756 0.9579 1.51121
TABLE XII: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{rp-gs}$ $mAUC_{svps}$
GAED 0.13484 13.79983
RGAE 0.12119 14.38906
Auto-Net 0.14676 52.05284
DFAN-HAD 0.35608 5.27753
DirectNet 0.12315 49.52139
BSDM 0.14287 8.40576
TGFA 0.13830 32.42845
DWSDiff 0.53694 67.20618
As can be seen in Table X, this table shows the accuracy of the anomaly detection methods, in which BSDM and DWSDiff are based on diffusion model, DWSDiff is in the first place in all other items except for the integrated detection performance (${mAUC}_{TD}$) which is second only to DFAN, while the data of BSDM is not outstanding (third in the integrated detection performance), but it shows the effect is not is not bad.
TABLE XIII: Comparison of Reasoning Time for Each Anomaly Detection Method
Methods Reasoning time (s)
Airport Beach Urban Average Time (s)
GAED 0.0300 0.0496 0.0257 0.0289
RGAE 0.0135 0.0251 0.0137 0.0141
Auto-Net 0.0020 0.0032 0.0020 0.0022
DFAN-HAD 0.0232 0.0448 0.0269 0.0264
DirectNet 0.0145 0.02839 0.0153 0.0145
BSDM 1.2306 1.2740 1.2509 1.2421
TGFA 1.6190 1.5505 1.5236 1.5392
DWSDUR 0.8016 0.8169 0.8192 0.8094
Table XI shows that both diffusion models have longer and slower inference times for each medium dataset scenario, with average times 99.8% higher (BSDM) and 99.7% higher (DWSDiff) than the shortest time (Auto-Net), which is hundreds of times higher than the optimal time, and is very slow in comparison. Again seeking performance DWSDiff is optimal and BSDM is a good choice.
Comparison of the experimental results of image recovery, as shown in Fig. 18, the overall effect of image recovery are relatively modest, but the methods based on the diffusion model (DDPM, EDiffSR, FastDiffSR) are better handled in terms of details, and the degree of recovery is closer to the high-resolution map:
Figure 18: Comparison of different datasets on hyperspectral image recovery methods: (a) High resolution map (b) HESNet [170] (c) TransENet [171] (d) NDSRGAN [172] (e) HAT [173] (f) DDPM [19] (g) EDiffSR [149] (h) FastDiffSR [148]
TABLE XIV: Comparison of Metrics for Different Image Recovery Methods on Potsdam Dataset
Methods Evaluation metrics(Potsdam)
PSNR FID LPIPS Params
HESNet 32.932 44.192 0.3014 5.4330
TransENet 32.768 41.310 0.2878 37.459
NDSRGAN 30.529 51.209 0.2137 17.510
HAT 32.987 44.932 0.2962 25.821
DDPM 31.989 24.307 0.0997 26.166
EDuFSR 30.860 36.669 0.1104 26.790
FastDuFSR 33.130 22.832 0.1152 23.229
As can be seen in Table XII, the three methods based on diffusion model (DDPM, EDiffSR, FastDiffSR) show unique advantages in key indicators, among which, the FsatDiffSR method has the best overall performance in terms of PSNR and FID, which is far superior to the traditional methods (compared to the second: (HAT) 0.43%, (DDPM) 6.07%, and (DDPM) 13.5% difference), indicating that its recovered image reaches a higher level in both detail fidelity and distributional consistency, but in terms of the number of parameters, the diffusion model-based methods are all higher than HESNet: 79.4%, 79.7%, 76.6%, respectively, thus affecting the training time too long.
TABLE XV: Comparison of Metrics for Different Image Recovery Methods on Toronto Dataset
Methods Evaluation metrics (Toronto)
PSNR FID LPIPS Params
HESNet 25.413 118.791 0.3234 5.5070
TransENet 25.565 97.958 0.3112 37.533
NDSRGAN 23.237 97.922 0.2534 17.510
HAT 25.368 119.05 0.2957 25.821
DDPM 24.733 66.583 0.1759 26.166
EDdfSR 23.990 81.412 0.1626 26.790
EasDdfSR 25.527 55.761 0.1561 23.229
Similar to Table XII, it can be seen from Table XIII that the three methods based on the diffusion model have a good overall performance, especially FastDiffSR has the best overall performance, and combining the two tables, the three diffusion methods have a higher quality of generation and more semantic information retention than the traditional methods. However, the three methods still have the problem of larger number of parameters in comparison.
TABLE XVI: Comparison of Measurement metrics for Different Image Recovery Methods
Methods Measurement metrics
Entropy NIQE
HESNet 7.0826 19.913
TransENet 7.0853 20.319
NDSRGAN 7.0528 19.214
HAT 7.1177 20.385
DDPM 7.1204 16.532
EDiIfSR 7.1838 17.641
EastDiffSR 7.1254 16.342
The entropy and generation quality metrics in Table XIV show that the three diffusion methods retain more semantic information and have higher generation and recovery quality, corresponding to Tables XII and XIII.Also combining the three tables, overall FastDiffSR has a balanced performance in terms of recovery accuracy and efficiency, DDPM is more advantageous in terms of perceptual quality, and EDiffSR can be used as a practical choice for moderate performance.
Comparison of noise suppression experimental results, as shown in Figure 19, with the noise map as well as the real noiseless map compared to the overall performance of the Diff-Unmix method has the best effect, followed by DDRM, and finally DDS2M are based on the diffusion model method, they are more successful noise suppression; in contrast to the traditional mainstream method effect is worse, to LLxRGTV in which the best:
Figure 19: Comparison of different datasets on hyperspectral noise suppression methods: (a) Noise map (b) True noise-free map (c) LLxRGTV [174] (d) FGSLR [175] (e) LRTDCTV [176] (f) DIP [177] (g) LLnRPnP [178] (h) HLRTF [110] (i) DDRM [179] (j)DDS2M [138] (k)Diff-Unmix [142]
TABLE XVII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,3))
PSNR SSIM FSIM SAM Time (s)
LLxRGTV 27.640 0.679 0.868 0.218 38
FGSLR 25.561 0.480 0.718 0.470 499
LRTDCTV 24.593 0.533 0.739 0.431 42
DIP 20.063 0.405 0.798 0.538 74
LLnRPhP 25.102 0.592 0.768 0.425 289
HLRTF 30.340 0.689 0.874 0.329 25
DDRM 27.810 0.786 0.893 0.387 23
DDS2M 30.078 0.666 0.834 0.355 318
Diff-Unmix 31.408 0.902 0.958 0.282 42
TABLE XVIII: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Gaussian noise: N(0,0,2))
PSNR SSIM FSIM SAM Time (s)
LLaRGUX 31.152 0.802 0.917 0.205 38
FGSLR 30.126 0.737 0.878 0.262 249
LRTDCTV 25.952 0.658 0.816 0.406 43
DIP 24.181 0.608 0.825 0.475 72
LLnRPhP 28.664 0.748 0.861 0.379 240
HLRTF 33.011 0.808 0.925 0.275 23
DDRM 29.412 0.865 0.922 0.293 20
DDS2M 32.804 0.786 0.895 0.334 354
Diff-Unmix 33.059 0.964 0.940 0.116 37
The last three methods (DDRM/DDS2M/Diff-Unmix) in the experiment are based on the diffusion model, as can be seen from Tables XV and XVI, the first choice of the KAIST two different noise situations in the dataset, these three methods in the noise suppression performance have their own characteristics, and the Diff-Unmix in the key indexes is the most outstanding performance, which PSRN, SSIM, and FSIM are the highest values, compared with other methods to show its significant advantages in noise suppression, structure preservation and spectral fidelity, but the time processing of the method under the diffusion model is slightly higher (HLRTF is the fastest), and the overall efficiency is still in the acceptable range.
TABLE XIX: Comparison of Metrics for Different Noise Suppression Methods on CAVE Dataset
Methods Evaluation metrics (CAVE)
PSNR SSIM FSIM SAM Time (s)
LLxRGTX 28.743 0.712 0.829 0.121 40
FGSLR 24.474 0.346 0.554 0.511 1758
LRTDCTV 24.394 0.444 0.631 0.315 47
DIP 20.480 0.554 0.864 0.313 66
LLnRPhP 24.809 0.445 0.668 0.284 246
HLRTF 30.308 0.770 0.861 0.185 133
DDRM 30.521 0.754 0.873 0.196 16
DDS2M 30.308 0.724 0.863 0.251 319
Diff-Unmix 32.714 0.940 0.957 0.129 43
Table XVII shows that the methods under the diffusion model are overall better than the other methods, of which Diff-Unmix remains the most prominent, demonstrating that the diffusion model method is overall superior to the other methods, with the best performance and the noise suppression presenting the optimal effect.
TABLE XX: Comparison of Metrics for Different Noise Suppression Methods on CAVE-Toy Dataset
Methods Evaluation metrics (CAVE-Toy)
PSNR SSIM FSIM SAM Time (s)
LLARGUX 26.450 0.752 0.874 0.118 39
FGSLR 21.525 0.573 0.798 0.241 1149
LRTDCTV 25.214 0.665 0.783 0.228 45
DIP 22.092 0.594 0.895 0.191 67
LLnRPhP 23.088 0.541 0.717 0.207 240
HLRTF 27.308 0.730 0.941 0.161 41
DDRM 27.886 0.858 0.910 0.159 16
DDS2M 29.344 0.844 0.977 0.111 320
Diff-Unmix 28.046 0.945 0.993 0.156 43
Table XVIII is roughly in agreement with Tables XV, XVI and XVII diffusion models method noise suppression performance is generally the best, and in combination with the above tables, the overall effect is highly acceptable, with obvious denoising effect, close to noiseless, and with better detail retention, despite the fact that the temporal processing is not particularly fast.
VI-B Key discoveries and Summaries from the Experiments
Through the experimental comparisons in Section V.B.2, the diffusion model shows significant advantages in hyperspectral image processing tasks, but also exposes some limitations.
In the classification task, the OA (93.15%, 94.78%, 98.97%), AA (96.44%, 93.84%, 99.47%) and Kappa coefficients of the SpectralDiff method based on the diffusion model on the Indian Pines, Pavia University and Salinas datasets ( 92.18%, 93.06%, 98.85%) are better than traditional deep learning methods (e.g., GAN, VAE), indicating that it can effectively capture the complex spectral-spatial features of hyperspectral data. However, its inference time (48.53 s, 256.23 s, 242.16 s) is significantly higher than that of lightweight models such as miniGCN (0.95 s, 2.40 s, 2.13 s), revealing the computational efficiency bottleneck of the diffusion model in high-dimensional data processing.
In the target detection task, the diffusion-based Diff-Mosaic method achieves optimal IoU (79.44%, 91.18%), detection probability (Pd: 99.99%, 99.47%), and false alarm rate (Fa: 3.19%, 1.91%) on the SIRST and NUDT-SIRST datasets, and is especially good at complex background suppression and target localization accuracy. However, the number of parameters (4.7M) and the inference time (43.42 s) are still higher than that of the traditional method ACM (0.29M, 18.53 s), which indicates that the diffusion model needs to be further lightweighted to meet the real-time demand.
In the anomaly detection task, DWSDiff comprehensively leads in mAUC metrics, such as mAUC(D,F) (0.97427), mAUC(BS) (0.95791), and mAUC(SNPR) (67.20618), which verifies its ability to suppress complex backgrounds through the dual-window strategy and iterative diffusion. However, its inference time (0.8094 s) is still much higher than that of Auto-AD (0.0022 s), and the model is sensitive to hyperparameters (e.g., noise scheduling), and its generalization ability needs to be further optimized, suggesting that the practical deployment of diffusion models in anomaly detection needs to balance accuracy and efficiency.
In image recovery tasks, diffusion model methods (e.g., FastDiffSR, EDiffSR, and DDPM) significantly outperform conventional methods in key metrics (PSNR, FID, LPIPS). For example, on the Potsdam dataset, FastDiffSR achieves a PSNR of 33.13 dB and a FID of 22.83, indicating that it outperforms conventional methods in both detail fidelity and distribution consistency. In addition, FastDiffSR achieves a PSNR of 25.53 dB and a FID of 55.76 on the Toronto dataset, which further validates its advantage in semantic information retention. However, the high number of parameters of the diffusion model method (e.g., 23.23M for FastDiffSR) leads to a significant increase in training time. It shows that the diffusion model has strong generative capability, but its computational efficiency and model complexity still need to be optimized to meet the actual deployment requirements.
In noise suppression tasks, diffusion model-based methods (e.g., Diff-Unmix, DDRM, and DDS2M) show significant advantages in metrics such as PSNR, SSIM, and FSIM. For example, on the KAIST dataset, Diff-Unmix achieves a PSNR of 31.41 dB, an SSIM of 0.902, and an FSIM of 0.958, indicating significant advantages in noise suppression, structure preservation, and spectral fidelity. In addition, on the CAVE dataset, Diff-Unmix has a PSNR of 32.71 dB and an SSIM of 0.940, further validating its superiority. However, the long inference time of the diffusion model method (e.g., 42 s for Diff-Unmix) restricts its application in scenarios with high real-time requirements, suggesting that the diffusion model is able to remove the noise and retain the details efficiently, but its computational efficiency still needs to be further improved.
Some key points are also found.1. Accuracy advantage: diffusion models are significantly mainstream deep learning methods (e.g., GAN, VAE) in tasks such as classification, target detection, and anomaly detection, especially in complex background suppression, high-dimensional feature modeling, and noise robustness.2. Computational cost: the computational complexity and inference time of diffusion models are generally high, e.g., SpectralDiff The inference time on Salinas dataset reaches 242 seconds, which is far more than miniGCN (2.13 seconds), limiting its application in real-time scenarios (e.g., disaster monitoring).3. Model adaptation: diffusion models are sensitive to hyper-parameters (e.g., noise scheduling, number of diffusion steps), and need to be customized with optimization strategies for hyperspectral data characteristics. For example, DWSDiff mitigates the background estimation bias by a double-window strategy, but its generalization still needs to be verified.4. Complementarity with traditional methods: traditional methods (e.g., SVM) are still advantageous in terms of interpretability and lightness, while diffusion models are more suitable for high-precision and complex scene tasks. Future research can explore hybrid architectures to balance efficiency and performance.
VII Outlook
In recent years, diffusion models have made significant progress in tasks such as image generation, denoising and reconstruction by virtue of their powerful generative capabilities and modeling advantages over complex data distributions. The diffusion model shows unique potential in the multidimensional and high-information-density field of hyperspectral image processing. Hyperspectral images contain spectral information in hundreds of consecutive bands, but their high-dimensional characteristics and noise interference pose challenges for traditional methods. As shown in the previous chapter, diffusion models can accurately model spectral-spatial features through a step-by-step denoising process, thus achieving higher accuracy in the task. However, the fact is that diffusion models still have a long way to go in the hyperspectral domain, and there are many tasks that still need to be explored and further improved, and the current research is still limited by the consumption of computational resources, insufficient cross-modal fusion and lack of interpretability. In the future, technical innovations can be promoted in the following three aspects: innovation in fusion with other traditional deep learning models, optimization and deployment in real scenarios, and automated model design and hyperparameter tuning.
1, Fusion innovation with other deep learning models, so that the diffusion model and the mainstream architecture to form a complementary, such as CNN + diffusion model, the use of CNN to extract spatial-spectral features as the a priori conditions of the diffusion process, to improve the efficiency of the generation; Transformer + diffusion model, through the attention mechanism to model the long-range dependence, to optimize the super-resolution reconstruction of hyperspectral images; GAN / VAE + diffusion model, which first generates low-resolution samples with GAN or VAE, and then performs detail enhancement and denoising by diffusion model to balance the generation speed and quality; Physics-driven fusion: combining physical a priori such as radiative transfer model, to enhance the generalization ability of diffusion model in complex environments (e.g., atmospheric interference).
2. Optimization and deployment in real-world scenarios need to be achieved through model lightweighting (e.g., knowledge distillation, dynamic diffusion step size), hardware adaptation (parallelized FPGA/GPU-based design) and embedded deployment (low-precision inference on edge devices). For vertical domains such as agricultural disease monitoring and mineral exploration, task-oriented diffusion frameworks need to be developed to balance accuracy and real-time requirements.
3. Automated model design and hyperparameter tuning. The performance of diffusion models is highly dependent on noise scheduling, network architecture and training strategies. Automated techniques (e.g., neural architecture search, meta-learning) can adaptively optimize the model architecture; reinforcement learning-driven hyperparameter search can dynamically adjust the denoising process for different hyperspectral tasks. Meanwhile, the development of unified evaluation benchmarks (e.g., cross-dataset generalization metrics) will accelerate the application and iteration of automated techniques.
In conclusion, diffusion model provides a new paradigm for hyperspectral image processing, but its development still needs to break through the bottlenecks of computational efficiency, cross-modal fusion and domain adaptation. In the future, through the combination of algorithmic innovation, hardware synergy and automation technology, the diffusion model is expected to realize smarter and more efficient analysis systems in the fields of environmental monitoring, precision agriculture, etc., and to promote hyperspectral technology from the laboratory to large-scale applications.
VIII Conclusion
In summary, compared with other deep generative models, the diffusion model is robust to the inherent noise in hyperspectral images, better adapts to its variability and complexity, and provides a more stable training process. Meanwhile, diffusion model with progressive denoising mechanism with high-dimensional data modeling capability, the diffusion model has injected a revolutionary impetus into the field of hyperspectral image analysis. It effectively solves the limitations of traditional methods in high-dimensional data processing, noise interference and small sample scenarios through the Markov chain noise injection and inverse generation process. Therefore, it is an inevitable trend for diffusion model to be applied in the field of hyperspectral images. This paper firstly introduces the related knowledge of hyperspectral as well as the basic theory of diffusion model, and then, reviews the research of diffusion model in the field of hyperspectral images, including: data enhancement, generation and complementation, classification segmentation, anomaly detection, target detection, denoising and other tasks, moreover, this paper fully demonstrates the superiority of diffusion model in this field by comparing some visualization results and quantitative indexes on hyperspectral image processing between diffusion model and traditional model. The limitations of the existing diffusion models in the hyperspectral domain are also discussed, and it is emphasized that further exploratory studies can be carried out for extended applications and model deployment.