### v2

Abstract
Hyperspectral image processing and analysis has important application value in remote sensing, agriculture and environmental monitoring, but its high dimensionality, data redundancy and noise interference etc. bring great challenges to the analysis. Traditional models have limitations in dealing with these complex data, and it is difficult to meet the increasing demand for analysis. In recent years, Diffusion models, as a class of emerging generative approaches, have demonstrated promising capabilities in hyperspectral image (HSI) processing tasks. By simulating the diffusion process of data in time, the Diffusion Model are capable of modeling high-dimensional spectral structures, generate high-quality samples, and achieve competitive performance in spectral-spatial denoising tasks and data enhancement. In this paper, we review the recent research advances in diffusion modeling for hyperspectral image processing and analysis, and discuss its applications in tasks such as high-dimensional data processing, noise removal, classification, and anomaly detection. The performance of diffusion-based models on image processing is compared and the challenges are summarized. It is shown that the diffusion model can significantly improve the accuracy and efficiency of hyperspectral image analysis, providing a new direction for future research.
keywords:
Hyperspectral image, Diffusion model, High dimensional data processing, Noise removal, Data enhancement, Anomaly detection.
1 Introduction
Hyperspectral image processing and analysis are crucial in remote sensing, agriculture, and environmental monitoring, offering significant application value. Continuous advancements in technology have led to improved resolution and broader spectral ranges in hyperspectral images, enabling more precise feature identification and classification. However, the inherent challenges of high dimensionality, data redundancy, and noise interference in these images often limit traditional processing methods. Consequently, the development of novel models and methods has become a key research focus. Recently, Diffusion Models (DMs), as an emerging generative model, have demonstrated unique advantages in addressing these complexities within hyperspectral image processing and analysis.
Hyperspectral imagery, with its ability to capture detailed spectral information across numerous narrow bands, offers a rich source of both spectral and spatial data. This high-dimensional and intricate data has found extensive applications in fields like agriculture, environmental monitoring, and mineral exploration. However, analyzing hyperspectral images presents several challenges. The sheer volume of high-dimensional data demands significant processing and storage capabilities, leading to high computational complexity. Moreover, the strong correlation among bands introduces data redundancy, complicating analysis. Issues such as noise interference and mixed pixel problems further impact the accuracy of the analysis.
Neural networks have dominated the past decade, profoundly impacting various domains, including image [1, 2], audio [3, 4], text [5], and point cloud processing [6], particularly with the rise of generative models. From a probabilistic standpoint, a key feature of a generative model is its ability to produce samples ($\widetilde{{x}}\sim p_{\theta}(\widetilde{{x}})$) that match the distribution of the training data (${x}\sim p_{\theta}({x})$). While Energy-Based Models (EBMs) achieve this by defining non-normalized probability densities, they require slow, iterative Markov Chain Monte Carlo (MCMC) sampling during both training and inference [7]. Recent advances in deep learning architectures have reignited interest in generative modeling, leading to improved visual fidelity and faster sampling speeds. This has seen the emergence of models such as Generative Adversarial Networks (GANs) [8], Variational Auto-Encoders (VAEs) [9], and Normalizing Flows [10]. In addition to these, Diffusion Models (DMs) offer a compelling alternative to existing VAEs, EBMs, GANs, and Normalizing Flows, as they don’t require aligning posterior distributions, estimating complex partition functions, introducing extra discriminator networks, or imposing network constraints. To date, diffusion models have proven useful in a wide range of applications, from generative tasks like image generation [11], image super-resolution [12], and image restoration [13], to discriminative tasks such as image segmentation [14], classification [15], and anomaly detection [16]. Recently, the hyperspectral community has seen an exponential increase in diffusion-based techniques, with numerous studies dedicated to applying diffusion models in various hyperspectral scenarios. Given the significant attention diffusion modeling has garnered from the research community, there’s a large body of literature on this topic. Consequently, this survey aims to provide a comprehensive overview of the latest advancements and applications of such models in the field of hyperspectral spectroscopy.
This paper differentiates itself from [17] in several key aspects. First, while [17] provides a broader overview of diffusion models in generalized remote sensing, encompassing multispectral, SAR, and panchromatic data, it lacks an in-depth analysis of the specific properties and challenges of hyperspectral imagery. Its focus is more on traditional tasks like image generation and super-resolution, aiming to establish a foundational mapping of diffusion modeling in remote sensing. In contrast, this paper specifically targets the hyperspectral modality, offering a more focused and detailed examination. Second, although [17] touches upon conditional diffusion models [11] and latent diffusion models [18], its exploration of emerging architectures stops at early applications before 2023. It also doesn’t systematically analyze practical engineering constraints or deeply discuss the unique technological challenges inherent in hyperspectral imagery.
Conversely, this paper provides a multi-perspective classification of diffusion models within the hyperspectral domain, specifically addressing core characteristics like high data dimensionality and spectral redundancy. This offers a systematic taxonomy for studying diffusion models and their applications in this field. Furthermore, we categorize existing diffusion models into Denoising Diffusion Probabilistic Models (DDPM) [19], Score-Based Generative Models (SGM) [20], and Stochastic Differential Equation-based Models (SDE) [21], analyzing their specific uses in the hyperspectral domain. We also systematically outline the progress of diffusion model applications across various tasks, offering a more comprehensive scope than [17]. These tasks are categorized into eight areas: data enhancement and synthesis, data generation and complementation, classification, segmentation, image generation, anomaly detection, noise suppression, and data recovery.
The significance of this paper is multifaceted: it fills a gap in the systematic review of diffusion models within the hyperspectral field; it provides a more comprehensive elaboration on the methodological innovations for adapting diffusion models to hyperspectral data; it uncovers unique challenges and optimization pathways for diffusion models in hyperspectral scenarios; and ultimately, it achieves a significant breakthrough across theoretical depth, technological breadth, and application value. This work lays a crucial theoretical and practical foundation for transitioning hyperspectral image processing from laboratory research to operational applications.
The paper is structured as follows: Section I defines hyperspectral images, outlining their characteristics, data properties, and primary analysis tasks. Section II delves into the theory, evolution, and variants of Diffusion Models, comparing them with traditional models and exploring their application in image generation. Section III focuses on the specific use of Diffusion Models in hyperspectral image analysis, detailing various applications. Section IV addresses the challenges and limitations inherent in Diffusion Models. Section V presents a comparative analysis between Diffusion Models and traditional methods, supported by experimental results. Section VI offers a future outlook on the role of Diffusion Models in hyperspectral imaging, and Section VII concludes the paper with a summary of its content.
2 Basics of hyperspectral imagery
2.1 Definition and Characteristics of Hyperspectral Imagery
A hyperspectral image is a spectral image characterized by a spectral resolution on the order of ${10}^{-2}\lambda$, encompassing tens to hundreds of continuously subdivided spectral bands across the visible to infrared light range. Captured by hyperspectral sensors (imaging spectrometers) mounted on various platforms, these images can cover the ultraviolet, visible, near-infrared, and mid-infrared regions of the electromagnetic spectrum within a target area. Crucially, each pixel in a hyperspectral image contains information from numerous spectral bands, forming a continuous spectral curve. This unique capability allows for the simultaneous acquisition of both spatial and spectral information, effectively integrating spectra with images.
Thanks to the maturation of imaging spectroscopy technology, hyperspectral images offer a unique fusion of image and spectral information. This integration provides a rich array of features, proving highly beneficial for both feature identification and processing tasks.
Image information captures external qualities like size, shape, and defects, while spectral information reveals internal physical structure and chemical composition. This dual capability allows for the acquisition of both multi-segment remote sensing imagery and continuous spectral data for each image element. This enables a multi-dimensional representation of feature characteristics, combining spatial and spectral information for more accurate identification and classification.
Hyperspectral images boast high spectral resolution, typically at the nanometer level. This allows for a detailed portrayal of feature reflectance spectra, enabling the differentiation of spectrally similar substances and significantly enhancing the accuracy of feature classification and identification. Simultaneously, these images contain an enormous amount of data, often comprising dozens to hundreds of bands, leading to information redundancy and multi-segmentation. This vast data volume necessitates high demands for storage and processing capacity. Despite these challenges, hyperspectral images provide exceptionally rich spectral information, capable of capturing subtle spectral differences and revealing correlations across both spatial and spectral dimensions.
Finally, hyperspectral images are characterized by their fine spectral feature expression, nonlinear characteristics, low signal-to-noise ratio, and high data dimensionality. These images are broadly applied across numerous fields, including agriculture, environmental monitoring, mineral exploration, and medicine, enabling the recognition and classification of diverse substances.
Figure 1: Hyperspectral image (a) and Illustration of hyperspectral image (b)
2.2 High-dimensional characterization of hyperspectral image data
Hyperspectral image data is inherently high-dimensional, primarily due to its multidimensional structure, significant data redundancy, high computational complexity, and exceptional information richness.
A hyperspectral image functions as a three-dimensional data cube, encompassing spatial, radiometric, and spectral dimensions. It features a large number of contiguous bands, forming high-dimensional vectors. The narrow bandwidths and nanometer-level resolution of these bands precisely reflect physicochemical properties. However, the high similarity between adjacent bands leads to information redundancy, increasing data complexity.
The "map in one" nature of hyperspectral images, combining spatial and spectral information, inherently elevates the data’s dimensionality. This necessitates considering both spatial and spectral characteristics during analysis, thereby increasing model complexity. Processing such high-dimensional data, especially for tasks like feature extraction, classification, and segmentation, demands substantial computational resources, leading to significantly increased computational complexity. Storage requirements are also high due to the immense data volume, which in turn incurs high costs. Traditional machine learning methods struggle with slow training in high-dimensional spaces and are susceptible to noise interference, a problem exacerbated in high dimensions, making processing more difficult. While deep learning can automate feature extraction, it often requires a large number of costly labeled samples.
The high dimensionality of hyperspectral data can lead to excessive resource waste, often resulting in the "curse of dimensionality" and diminished performance. To counter this, dimensionality reduction techniques like PCA and LDA are crucial for preserving essential information while simplifying the model. During this complex process, it is vital to mitigate issues such as overfitting during training.
2.3 Main tasks in hyperspectral image analysis
The primary tasks in hyperspectral image analysis, including classification, segmentation, and anomaly detection, hold significant value in processing these complex, high-dimensional datasets. They are essential for extracting meaningful and useful information. Let’s delve into each of these tasks:
Classification in hyperspectral imagery involves assigning each pixel or region to a predefined class, such as vegetation, water bodies, or buildings. This is widely applied in agriculture for crop classification and pest detection, in environmental monitoring for land use classification and forest cover monitoring, and in mineral exploration for identifying mineral types. However, several challenges persist: the high-dimensional nature of hyperspectral data increases classification complexity; data labeling is costly due to the extensive amount of labeled data required; and the mixed pixel problem, where a single pixel encompasses multiple features due to spatial resolution limitations, complicates accurate classification. Traditional approaches include statistical methods (e.g., maximum likelihood classification) and machine learning methods (e.g., Support Vector Machines, SVM). More recently, deep learning methods such as Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs) have also been employed.
Segmentation in hyperspectral imagery involves partitioning an image into distinct regions or objects that share similar spectral characteristics. This can be performed at either the pixel or object level. Common applications include segmenting tumor regions in medical image analysis, identifying buildings and roads in urban planning, and delineating crop areas in agriculture. However, several challenges arise: the high dimensionality of hyperspectral data complicates the segmentation process; noise interference can compromise accuracy; and the mixed pixel problem, where a single pixel represents multiple features, makes precise segmentation difficult. Traditional approaches often rely on clustering-based methods (e.g., K-means) or edge detection techniques. Deep learning methods, such as Fully Convolutional Networks (FCNs) and UNet, are increasingly employed for this task.
Target detection in hyperspectral imagery involves identifying and locating specific targets or objects, such as military assets, mineral resources, or vegetation types, within an image. Its applications are diverse, ranging from military target identification and tracking, to the detection of specific minerals in exploration, and the identification of particular vegetation or pollution sources in environmental monitoring. This task presents several challenges: the high dimensionality of hyperspectral data increases the complexity of detection; target diversity, where different targets may exhibit similar spectral characteristics, can hinder accurate detection; and background interference from complex environments can significantly impact detection accuracy. Traditional methods for target detection include spectral matching techniques and statistical methods. More recently, deep learning approaches like Region-based Convolutional Neural Networks (RCNN) and YOLO (You Only Look Once) have been widely adopted.
Anomaly detection in hyperspectral imagery identifies pixels or regions that significantly deviate from their surrounding environment. These anomalies can represent unknown targets or unique features. This technique finds applications in military contexts for detecting unidentified targets, in environmental monitoring for pinpointing pollution sources, and in agriculture for the early detection of pests and diseases. However, challenges arise due to the high dimensionality of hyperspectral images, which increases the complexity of anomaly detection. Furthermore, the diversity of anomalies, where different types may exhibit varying spectral characteristics, complicates their detection. Lastly, background interference from complex environments can negatively impact detection accuracy. Traditional approaches include statistical methods (e.g., the RX algorithm) and clustering-based methods. More recently, deep learning methods such as Autoencoders and Generative Adversarial Networks (GANs) have been employed.
Noise suppression in hyperspectral imagery aims to reduce or eliminate unwanted noise through specialized algorithms and techniques. This process is crucial for enhancing image quality and ensuring the accuracy of subsequent analyses. Its applications span diverse fields, including agriculture, military operations, the food industry, and mining exploration. Common methods for noise suppression primarily involve frequency or transform domain noise reduction, spatial noise reduction, and deep learning-based noise reduction.
Hyperspectral data recovery is a more advanced form of noise suppression, focused on removing noise and interference from hyperspectral images to restore their original spectral and spatial information. This process significantly enhances data quality and usability. Key application areas include agriculture, environmental monitoring, biomedicine, and geological exploration. Common methods for data recovery include spectral reconstruction, spectral dimension recovery, low-rank representation modeling, and various deep learning approaches.
Hyperspectral image data augmentation and synthesis involves generating new hyperspectral image samples to increase the diversity and quantity of training data, thereby enhancing a model’s generalization capability. This is frequently applied in model training to diversify data and improve model robustness, and in data completion to generate missing hyperspectral image data.Challenges in this area include: the high dimensionality of hyperspectral images, which complicates data augmentation and synthesis; and ensuring the quality and diversity of generated samples to maintain model validity. Traditional methods often rely on interpolation-based techniques and statistics-based methods. More recently, deep learning approaches like Generative Adversarial Networks (GANs) and Diffusion Models have shown promise.
Hyperspectral data generation and completion is a fundamental task in data science that addresses issues like missing data, noise interference, and data scarcity, while also expanding application possibilities through synthetic data. Data generation involves creating synthetic data that aligns with real data distributions, often to augment dataset scale or simulate specific scenarios. Data completion, on the other hand, focuses on filling in missing values or repairing incomplete data to ensure consistency and completeness. These techniques are commonly employed in fields such as remote sensing, medical imaging, and agricultural monitoring. Traditional methods include statistics-based approaches (e.g., interpolation) and machine learning techniques (e.g., K-nearest neighbor, decision trees). More recently, deep learning methods like GANs, VAEs, and Diffusion Models have become prominent for these tasks.
In summary, the core tasks of hyperspectral image analysis encompass classification, segmentation, target detection, anomaly detection, and data enhancement and synthesis. These tasks are vital in processing hyperspectral imagery, enabling the extraction of valuable information from complex, high-dimensional datasets. While these tasks contend with challenges such as high-dimensional data processing, noise interference, and mixed pixel issues, combining traditional and deep learning methodologies can significantly enhance the accuracy and efficiency of hyperspectral image analysis.
3 Diffusion Model
The Diffusion Model (DM) is an innovative generative model that has recently showcased remarkable capabilities in image generation, denoising, and data enhancement. Its core principle involves modeling a data diffusion process over time, progressively transforming noisy data into desired target data. Specifically, DMs operate by gradually converting original data into noise through a forward diffusion process, and then meticulously reconstructing the original data from this noise via a reverse diffusion process. This methodology proves highly effective for handling high-dimensional data, as well as for image generation and denoising tasks, demonstrating particular advantages in the realm of hyperspectral image processing.
3.1 Theoretical Foundations of the Diffusion Model
The Diffusion Model operates through two distinct processes: forward diffusion and backward diffusion. The training objective is to minimize the discrepancy between these two processes. Fundamentally, diffusion models are typically categorized into three main types: (1) Denoising Diffusion Probabilistic Models (DDPM), (2) Score-Based Generative Models (SGM), and (3) Stochastic Differential Equation-based Models (SDE/NCSN). The foundational principles of each of these three diffusion models are described below.
3.1.1 Denoising Diffusion Probabilistic Model (DDPM)
Denoising Diffusion Probabilistic Models (DDPMs) [19, 22] employ two Markov chains. A forward chain systematically introduces noise to transform the data into a simple prior distribution, typically a standard Gaussian. Conversely, an inverse chain learns to reverse this process, converting the noise back into data through a transition kernel parameterized by a deep neural network.
New data points are generated by first randomly sampling a vector from the prior distribution, then using an inverse Markov chain to ancestrally sample it [23].
Figure 2: Diffusion processes in diffusion models [19]
Formally, given a data distribution $\mathbf{x}_{0}\sim q(\mathbf{x}_{0})$, a forward Markov process generates a sequence of random variables, $\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}$, through a transition kernel $q(x_{t}\mid{x_{t-1}})$. By applying the probability chain rule and the Markov property, the joint distribution of these variables, conditional on $\mathbf{x}_{0}$, is then decomposed and denoted as $q(x_{1},\ldots,x_{T}\mid{x_{0}})$, which results in:
$q(\mathbf{x}_{1},...,\mathbf{x}_{T}\mid\mathbf{x}_{0})=\prod_{t=1}^{T}q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1}).$ (1)
In DDPMs, the transition kernel $q(x_{t}\mid{x_{t-1}})$ is specifically designed to gradually transform the data distribution $q(\mathbf{x}_{0})$ into a manageable prior distribution. A common and effective choice for this transition kernel is a Gaussian perturbation, typically defined as:
$q(\mathrm{x}_{t}\mid\mathrm{x}_{t-1})=N(\mathrm{x}_{t};\sqrt{1-\beta_{t}}\mathrm{x}_{t-1},\beta_{t}\mathrm{I}).$ (2)
where $\beta_{t}\in(0,1)$ is a hyperparameter set before model training. While other kernel types exist, this Gaussian transition kernel is used to simplify the discussion. As noted by Sohl-Dickstein et al. (2015) [22], this specific kernel allows for the marginalization of the joint distribution in equation (1), yielding an analytical form for $q(x_{t}\mid\mathbf{x}_{0})$ across all $t\in\{0,1,\cdots,T\}$. Specifically, with $\alpha_{t}:=1-\beta_{t}$ and $\bar{\alpha}_{t}{:}=\prod_{s=0}^{t}\alpha_{s}$:
$q(\mathbf{x}_{t}\mid\mathbf{x}_{0})=\mathcal{N}\left(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I}\right).$ (3)
Intuitively, this forward process gradually introduces noise into the data until its original structure is completely lost. To generate new data samples, the DDPM first samples an unstructured noise vector from the prior distribution (which is typically straightforward to obtain). It then progressively removes this noise by running a learnable Markov chain in the reverse temporal direction. Specifically, this inverse Markov chain is parameterized by the prior distribution $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ and the learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$. We choose $p(\mathbf{x}_{T})=\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$ for the prior because the forward process is designed such that $q(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{x}_{T};0,\mathbf{I})$. The learnable transition kernel $p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})$ employs:
$p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\right).$ (4)
where $\theta$ represents the model’s parameters, while the mean $\mu_{\theta}(\mathbf{x}_{t},t)$ and variance $\Sigma_{\theta}(\mathbf{x}_{t},t)$ are parameterized by a deep neural network.
The key to successfully training the network lies in ensuring the reverse process $p_{\theta}(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$ accurately matches the true time reversal of the forward process $q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})$. This is typically achieved by minimizing the Kullback-Leibler (KL) divergence between the two:
$\mathrm{KL}\left(q(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})|\||p_{\theta}(\mathrm{x}_{0},\mathrm{x}_{1},\cdots,\mathrm{x}_{T})\right)$ (5)
=$\overset{(i)}{\operatorname*{=}}-\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})}[\log p_{\theta}\left(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T}\right)]+\mathrm{const}$ (6)
=$\begin{split}\overset{(ii)}{\operatorname*{=}}\mathbb{E}_{q(\mathbf{x}_{0},\mathbf{x}_{1},\cdots,\mathbf{x}_{T})}\left[-\mathrm{log}p(\mathbf{x}_{T})-\sum_{t=1}^{T}\log\frac{p_{\theta}(\mathbf{x}_{t-1}\mid\mathbf{x}_{t})}{q(\mathbf{x}_{t}\mid\mathbf{x}_{t-1})}\right]\\ +\mathrm{const}:=-L_{NLB}(\mathbf{x}_{0}).\end{split}$ (7)
$\overset{(iii)}{\operatorname*{\operatorname*{\geq}}}\mathbb{E}[-\log p_{\theta}\left(\mathbf{x}_{0}\right)]+\mathrm{const}$ (8)
where "const" is a constant independent of the model parameters $\theta$. The objective of DDPM training is to maximize the variational lower bound, or equivalently, to minimize the negative variational lower bound. This can be efficiently estimated through Monte Carlo sampling [24] and optimized effectively via stochastic optimization [25].
3.1.2 Score-Based Generative Models (SGM/NCSN)
At the core of Score-Based Generative Models (SGM/NCSN) [26, 20, 27] lies the concept of a (Stein) score, also known as a fraction or score function [28]. For a given probability density function $p(\mathbf{x})$, the score function is defined as the gradient of the logarithmic probability density: $\nabla_{\mathbf{x}}\log p(\mathbf{x})$. Unlike the Fisher score, $\nabla_{\theta}\log p_{\theta}(\mathbf{x})$, which is commonly used in statistics and depends on the model parameters $\theta$, the Stein score is a function of the data $\mathbf{x}$ itself. Essentially, it’s a vector field that points in the direction of the steepest ascent of the probability density function.
The central idea behind Score-Based Generative Models (SGMs) [20] is to progressively add Gaussian noise to data and then collectively estimate the score function for all these noisy data distributions. This estimation is achieved by training a deep neural network, specifically known as a Noise Conditioned Scoring Network (NCSN) [20], which is conditioned on the noise level. Samples are then generated by linking these score functions across decreasing noise levels. This is done using various score-based sampling methods, including Ronzmann Monte Carlo [29, 30, 31, 20, 21], stochastic differential equations [32, 21], ordinary differential equations [33, 34, 35, 36, 21], and their combinations [21]. A key advantage of the SGM formulation is the complete decoupling of training and sampling, which allows for the use of multiple sampling techniques once the score function has been estimated.
[…]
$\mathbb{E}_{t\sim\mathcal{U}[[1,T]],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[\lambda(t)\|\epsilon-\epsilon_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].$ (9)
$\begin{split}=\mathbb{E}_{t\sim\mathcal{U}\|1,T\|\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\sigma_{t}^{2}\|\nabla_{\mathbf{x}_{t}}\log q\left(\mathbf{x}_{t}\right)-\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\|^{2}\right].\end{split}$ (10)
=$\begin{split}\overset{(i)}{\operatorname*{=}}\mathbb{E}_{t\sim u[[1,T]]\mathbf{x}_{0}\sim q(\mathbf{x}_{0})\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}[\lambda(t)\sigma_{t}^{2}\\ \prod\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})-\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\parallel^{2}]+\mathrm{const}.\end{split}$ (11)
=$\begin{split}\overset{(ii)}{\operatorname*{=}}\mathbb{E}_{t\sim U[1,T],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\left\|-\frac{\mathbf{x}_{t}-\mathbf{x}_{0}}{\sigma_{t}}-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\right\|^{2}\right]+\mathrm{const}.\end{split}$ (12)
$\begin{split}(iii)\\ =&\mathbb{E}_{t\sim\mathcal{U}[[1,T]],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\epsilon\sim\mathcal{N}(0,\mathbf{I})}[\lambda(t)\|\epsilon+\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x}_{t},t)\|^{2}]\\ +\mathrm{const}.\end{split}$ (13)
where $\lambda(t)$ represents a positive weighting function, while $\mathbf{x_{t}}$ is derived from $\varepsilon$ and $\mathbf{x}_{0}$. $\mathcal{U}[[1,T]]$ denotes a uniform distribution over the set $\{1,2,\ldots,T\}$, and $\varepsilon$ is a deep neural network with parameters $\theta$, tasked with predicting the noise vectors for a given $\mathbf{x}_{t}$ and $t$. Comparing the equations, the training objectives for DDPM and SGM become equivalent once $\epsilon_{\theta}(\mathbf{x},t)=-\sigma_{t}\mathbf{s}_{\theta}(\mathbf{x},t)$ is established.
3.1.3 Diffusion model based on stochastic differential equations (Score-SDE)
Both DDPM and SGM can be further generalized to an infinite time step or noise level. In this advanced formulation, known as Score SDE [21], the perturbation and denoising processes are governed by the solution of a stochastic differential equation (SDE). This approach is employed for both noise perturbation and sample generation, where the denoising process necessitates estimating a score function of the noisy data distribution.
The Stochastic Differential Equation (SDE) was employed to score the data using a diffusion process, which is governed by the following SDE [21]:
$\mathrm{dx}=\mathrm{f}(\mathrm{x},t)\mathrm{d}t+g(t)\mathrm{dw}.$ (14)
where $f(\mathbf{x},t)$ and $g(t)$ represent the drift and diffusion functions of the SDE, respectively, while $w$ denotes a standard Wiener process (also known as Brownian motion). The forward processes in both the DDPM and SGM are essentially discretizations of this SDE. As demonstrated by Song et al. in 2020 [21], the corresponding SDE for the DDPM is given by $dx=-\frac{1}{2}\beta(t)xdt+\sqrt{\beta(t)}dw$. For the SGM, the corresponding SDE is $\mathrm{dx}=\sqrt{\frac{\mathrm{d}[\sigma(t)^{2}]}{\mathrm{d}t}}\mathrm{dw}$. This holds true as $T$ approaches infinity, where $e\beta(\frac{t}{T})=\beta_{t}T$ and $\sigma(\frac{t}{T})=\sigma_{t}$.
Here, $q_{t}({x})$ represents the distribution of $\mathbf{x}_{t}$ in the forward process. For any diffusion process structured like equation (14), Anderson [42] demonstrated that it can be inverted by solving the following reverse-time SDE:
$\mathrm{d}\mathbf{x}=[\mathrm{f}(\mathbf{x},t)-g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})]\mathrm{d}t+g(t)\mathrm{d}\overline{\mathbf{w}}.$ (15)
where $\bar{\mathbf{W}}$ represents the standard Wiener process as time flows backward, and $dt$ signifies an infinitesimal negative time step. The solution trajectories of this reverse SDE share the same marginal density as those of the forward SDE, but evolve in the opposite temporal direction [21]. Conceptually, the solution of a reverse-time SDE describes a diffusion process that progressively transforms noise into data. Furthermore, Song et al. (2020) [21] demonstrated the existence of an ordinary differential equation (ODE), known as the probability flow ODE, whose trajectory exhibits the same marginals as the reverse-time SDE. The probability flow ODE is defined by the following equation:
$\mathrm{d}\mathbf{x}=\left[\mathrm{f}(\mathbf{x},t)-\frac{1}{2}g(t)^{2}\nabla_{\mathbf{x}}\mathrm{log}q_{t}(\mathbf{x})\right]\mathrm{d}t.$ (16)
Both the inverse time SDE and the probabilistic flow ODE allow for sampling from the same data distribution because their trajectories share identical marginals. If the fractional function $\nabla_{\mathbf{x}}\log q_{t}(\mathbf{x})$ is known for each time step $t$, its inverse-time equations can be derived. These equations can then be used to generate samples via numerical techniques such as annealed Ronzwan dynamics [20], numerical SDE solvers [32, 21], numerical ODE solvers [33, 34, 21, 36, 43], and predictor-corrector methods (a combination of MCMC and numerical ODE/SDE solvers) [21]. Similar to the SGM, the transient fractional model $\mathbf{s}_{\theta}(\mathbf{x}_{t},t)$ is parameterized, and the fractional function is estimated by generalizing the fractional matching objective in Equation (13) to continuous time, resulting in the following objective:
$\begin{split}\mathbb{E}_{t\sim\mathcal{U}[0,T],\mathbf{x}_{0}\sim q(\mathbf{x}_{0}),\mathbf{x}_{t}\sim q(\mathbf{X}_{t}|\mathbf{X}_{0})}\\ \left[\lambda(t)\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q_{0t}\left(\mathbf{x}_{t}\mid\mathbf{x}_{0}\right)\|^{2}\right].\end{split}$ (17)
where $\mathcal{U}[0,T]$ denotes the uniform distribution over the interval $[0,T]$.
The preceding discussion has introduced the fundamental principles and formulas of the basic diffusion model. This theoretical framework underpins the success of diffusion models in modern image processing, enabling them to overcome certain limitations of traditional models, albeit with their own inherent drawbacks. In the context of hyperspectral imaging, research primarily focuses on Denoising Diffusion Probabilistic Models (DDPMs), which are the most frequently applied diffusion models in this domain. While the other two types have seen comparatively less application, they will also be explored in subsequent sections.
3.2 Evolution of the Diffusion model with different variants
The genesis of Diffusion Models dates back to 2015, when Sohl-Dickstein et al. first introduced their theoretical framework, uniquely merging nonequilibrium thermodynamics with generative models [22]. In 2019, Song Yang et al. advanced the field with Score-Based Generative Models (SGM/NCSN) [20], which leverage score matching to generate data by estimating the logarithmic gradient of the data distribution. This was further refined in 2020 with NCSN++ (Noise-Conditioned Score Networks++) [27], optimizing the network architecture (through residual connections and adaptive normalization), noise scheduling, and introducing a more stable training objective. The same year saw Ho et al.’s seminal work on DDPM (Denoising Diffusion Probabilistic Models) [19], which, by utilizing fixed noise scheduling and reparameterization, established the foundation for modern diffusion models.
Addressing DDPM’s limitations in generation quality, training efficiency, and log-likelihood optimization, Nichol et al. introduced Improve-DDPM (IDDPM) [44] in 2021. This variant incorporated learnable variance, cosine noise scheduling, and a hybrid loss function, while also accelerating sampling through time-step-importance sampling. Despite these improvements, IDDPM still faced complexities in computation, dataset limitations, insufficient generalization, and trade-offs between generation quality and speed. Concurrently in 2021, Song et al. proposed DDIM (Denoising Diffusion Implicit Models) [35], accelerating sampling via non-Markovian chains. Significantly, Rombach et al. unveiled the Latent Diffusion Model (LDM) [18], which dramatically reduced computational cost and maintained generation quality by shifting the diffusion process to a latent space rather than operating directly on pixels. Further enhancements for NCSN++ emerged with Score-SDE (based on stochastic differential equations) [21] and ODE (probabilistic streaming) [21], both optimizing the sampling process through stochastic and constant differentiation, respectively. Also in 2021, Dhariwal and Nichol introduced Classifier Guidance (Conditional Diffusion Model/CDM) [11], significantly boosting generation quality by integrating a classifier to guide the process.In 2022, Rombach et al. released Stable Diffusion [18], an advancement based on the LDM model specifically enhanced for text-to-image generation.
Since then, numerous important variants have been developed [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57], though none have yet reached the "landmark" status of the aforementioned models. A timeline illustrating the evolution of the diffusion model and its principal variants is provided below.
Figure 3: Evolution timeline of the Diffusion model
3.3 Traditional Deep Learning Models
Generative Models (GMs) are a class of machine learning models capable of learning data distributions and generating novel samples. Their applications are widespread in fields like computer vision and natural language processing, encompassing tasks such as data generation, data enhancement, image restoration, and anomaly detection. In this paper, we will primarily focus on traditional generative models, specifically Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), and autoregressive models.
The Generative Adversarial Network (GAN), proposed by Ian Goodfellow et al. in 2014 [8], is another significant framework for unsupervised anomaly detection, rooted in the concept of a zero-sum game. At its core, a GAN generates data through the adversarial training of two neural networks: a generator and a discriminator. The generator’s role is to produce data, while the discriminator’s task is to differentiate between this generated data and real data. This adversarial process drives the generator to progressively learn how to create data that closely resembles the distribution of real data. While GANs offer fast generation, they often suffer from unstable training and are prone to issues like mode collapse [58, 59, 60].
The Variational Auto-Encoder (VAE) [9] is a generative model that blends the strengths of Autoencoders with probabilistic graphical models. Its core idea is to learn latent representations from data and subsequently generate new samples. VAEs achieve this by using Variational Inference (VI) to approximate the data’s latent distributions. While they offer stable training, the quality of the generated data can sometimes be lower [61, 62, 63].
Autoregressive Models (ARMs) [64] are a class of generative models that operate by assuming the current data point in a sequence is dependent on preceding ones. Their core idea is to leverage this temporal dependence to generate new data points sequentially. Examples include Recurrent Neural Networks (RNNs), which employ a recursive structure for data processing, and Long Short-Term Memory (LSTM) Neural Networks, a specialized type of RNN. While ARMs are known for generating high-quality data, their step-by-step generation process can result in slower speeds [65, 66].
3.4 Successful Application of Diffusion Modeling in Image Generation
Since 2020, Diffusion Models have seen widespread application in image processing, primarily leveraging Denoising Diffusion Probabilistic Models (DDPMs). Early works like the initial DDPM [19], IDDM [44], and DDIM [35] optimized generation quality through Markov chains and refined noise addition/denoising processes, predicting noise residuals. However, these models were characterized by numerous sampling steps and slow computation.
In 2021, the OpenAI team introduced ADM-G (Classifier-Guided-Diffusion) [11], which, by integrating classifier-guided generation, allowed the model to surpass GANs in FID metrics for the first time. Despite this breakthrough, it remained dependent on classifiers and incurred high computational costs. To enhance text-image alignment and super-resolution, the Google Research team later introduced Imagen [46], combining the T5-XXL text encoder with a cascaded diffusion architecture, resulting in outstanding text alignment capabilities but demanding substantial resources.
Seeking multimodal and efficient generation, Robin Rombach et al. proposed Stable Diffusion in 2023 [18]. This model significantly reduced computational complexity by compressing the diffusion process into a latent space, albeit with some information loss. Concurrently, the OpenAI team also released CLIP-Guided-Diffusion [49], which integrates the CLIP model to enable cross-modal generation and support complex semantic control, ensuring semantic consistency. However, this model also requires significant training resources and has a slower generation speed.
More recently, in 2024, the Wisdom Source Institute introduced OmniGen [53], the first unified framework supporting tasks like text generation and image editing without additional plug-ins. Following this, the ByteDance team launched D-Dit [55], which combines continuous and discrete diffusion techniques to achieve bidirectional text-image generation, support multimodal tasks, and improve parameter efficiency by approximately 20%. Looking ahead to 2025, Yifan Zhou et al. are set to propose AF-LDM [57], aiming to enhance video generation consistency and model robustness by introducing shift isovariance loss and improving the UNet attention module to reduce potential spatial noise bias.
Simultaneously, the diffusion model is poised for continued innovation and development within the hyperspectral domain. Future advancements in this area will build upon existing diffusion models, with a focus on resolving challenges in hyperspectral image classification, anomaly detection, and recovery.
Figure 4: Improved division of the diffusion model according to different variants
Table 1: Summary of basic diffusion models, variants and their characteristics
Basic Diffusion Model Major Variants Reasons for Improvement Improvement Point Drawbacks Applications
DDPM IDDPM Improve generation quality and training stability. Introduction of cosine noise scheduling and hybrid objective function. to optimize the training process. Generation is still slow. Image generation, audio synthesis.
DDIM Accelerating the generation process. Non-Markov chain inverse process allowing large step size sampling. Sacrificing some generative diversity for speed. Fast image generation, interpolation.
Classifier-Guided Enhanced generation control. Using classifier gradients to guide the direction of generation and improve the quality of conditional generation. Additional training of classifiers is required, increasing computational cost. Conditional image generation, text-to-image.
LDM/Stable Reduced computational costs. Diffusion in latent space to reduce computational overhead in high-resolution images. Latent space compression may lose detail. High-resolution image generation, art creation.
SGM/NSCN NSCN++ Enhancement of model presentation skills. Introducing deeper network structure and improved score matching goals. High training complexity. Complex distribution modeling, scientific simulation.
Score-SDE Harmonization of diffusion frameworks. Modeling Continuous Diffusion Processes by Stochastic Differential Equation (SDE) for More General Theory. The sampling process is computationally intensive. Multimodal generation, cross-domain generation.
ODE Improved generation process stability. Replacing SDE with ordinary differential equations (ODE) for deterministic sampling. No significant increase in generation speed compared to SDE. Controlled generation, inverse problem solving.
4 Applications in Hyperspectral Image Processing and Analysis
Deep learning models are widely adopted across various fields, demonstrating significant advantages. In recent years, scholars have successfully applied them to hyperspectral imaging, yielding numerous achievements. Their application in hyperspectral image analysis spans several areas, including data enhancement and synthesis, data generation and completion, image classification, segmentation, target detection, and anomaly detection. This section will categorize hyperspectral image processing into two main approaches: those using traditional deep learning models and those employing diffusion-based models. The following provides a detailed description of these methods.
Figure 5: Deep learning model-based hyperspectral image processing
Figure 6: Timeline of the birth of each model
4.1 The main Diffusion model used in the field of hyperspectral imagery
Diffusion Model based Hyperspectral Image Processing was first applied to the hyperspectral field in 2022. Research indicates that Denoising Diffusion Probabilistic Models (DDPM) are predominantly used in this domain. This section will specifically analyze the technical aspects of DDPM and its variants within hyperspectral image processing. The primary diffusion models leveraged for this purpose include DDPM [19], IDDPM [44], DDIM [35], LDM [18], Classifier-Guided Diffusion [11], and Stable Diffusion [18].
We’ve already covered the basic principles of DDPMs. Since the fundamental principles and formulas of IDDPM and DDIM are essentially identical to DDPM, we won’t reiterate them here. LDM (Latent Diffusion Model) and Stable-Diffusion fall under the same category, sharing core ideas, with Stable-Diffusion effectively being a specific implementation of LDM. Therefore, the following sections will introduce the basic principles and formulas of LDM and Classifier-Guided (Conditional Diffusion Model).
4.1.1 Conditional Diffusion Model (CDM)
Similar to the evolution of GANs, diffusion models initially focused on unconditional generation, swiftly transitioning to conditional generation [67]. While unconditional generation is often used to push the boundaries of a model’s capabilities, conditional generation holds more practical value for applications, as it enables the output to be controlled based on human intent.
The introduction of conditioning into diffusion models first appeared in [11]. This pioneering work guided the generation of diffusion models by incorporating classifiers into pre-trained models, leading to the term "guided diffusion models." While this method offered low training costs, it increased inference costs by using classification results to steer the diffusion model’s sampling process, and it struggled to precisely control details for optimal image output. Consequently, the Google team [68] opted for a simpler approach: retraining the DDPM conditionally to control generated results, a technique they named Classifier-Free Guidance.
Formally, when given conditional information $\mathbf{c}$, the combined DDPM distribution changes from equations (2) and (3). Consequently, both the optimization objective and the sampling process are modified to take the following form:
$\mathcal{L}_{\mathrm{con}}(\theta)=\mathbb{E}_{x_{0},\epsilon,c,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon,c,t\right)\right\|^{2}\right].$ (18)
$y_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(y_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\tilde{\alpha}_{t}}}\epsilon_{\theta}(y_{t},c,t)\right)+\hat{\beta}_{t}z.$ (19)
This approach has seen broader adoption than guided diffusion models and forms the foundation for many significant diffusion models, including DALL-E2 [45], Imagen [49], and Stable Diffusion (SD) [18].
4.1.2 Latent Diffusion Model (LDM)
Directly manipulating diffusion models in the original image pixel space incurs high computational costs. To address this, Rombach et al. [60] proposed training and performing inference with diffusion models within a low-dimensional latent space, terming this approach the Latent Diffusion Model (LDM).
Figure 7: Architecture of LDM, where E and D denote the encoder and decoder, respectively [18]
As depicted in Fig. 7, the Latent Diffusion Model (LDM) comprises three distinct parts. Firstly, the leftmost gray box illustrates the LDM’s specific diffusion process: it shifts its operational domain from pixel space to a low-dimensional latent space by utilizing a pre-trained autoencoder. Within this space, noise is incrementally added to feature $\mathbf{z}$, yielding $\mathbf{z_{T}}$ as part of the forward diffusion. Secondly, the central orange box, also in latent space, houses a U-Net with cross-attention that supports multimodal inputs. This U-Net facilitates the reverse diffusion, restoring $\mathbf{z_{T}}$ back to $\mathbf{z}$. Lastly, the rightmost box contains the conditional encoder, which transforms various conditions (such as text and images) into feature vectors $\tau_{\theta}$. These vectors are then fed into the U-Net to guide image generation. The LDM’s ability to effectively reduce computational cost without sacrificing image synthesis quality, coupled with its accelerated inference speed, has led to its widespread adoption in natural scene generation and offers significant inspiration for applying diffusion models to hyperspectral applications.
4.2 Application of Diffusion model in image processing
4.2.1 Data Enhancement and Synthesis
In data enhancement and synthesis, super-resolution techniques play a crucial role in improving image quality. Super-resolution aims to reconstruct high-resolution images from their low-resolution counterparts [69]. However, these images can be degraded by factors like imaging equipment, weather conditions, or downsampling. This loss of detail is often more pronounced in hyperspectral images compared to natural images, making the inverse problem of reconstructing high-resolution versions significantly more challenging.
Before the advent of diffusion models, approaches based on deep neural networks, such as Convolutional Neural Networks (CNNs) and Transformers, sought to integrate feature information in a single step by finding suitable mapping functions. Methods like self-supervised learning [70], adaptive marker selection [71], and frequency information assistance [72] were employed to achieve this. In contrast, diffusion model-based approaches fuse low-resolution images and bootstrap images at each step of the diffusion process. This iterative integration of different image information has proven highly effective and has garnered significant attention in the hyperspectral field.
Although Hyperspectral Images (HSIs) boast high spectral resolution, their spatial resolution is often comparatively low [73], which can limit the effectiveness of various HSI-based applications. To overcome this, two primary methods are employed to achieve high-resolution hyperspectral maps: panchromatic sharpening [74] and multispectral and HSI fusion [75].
Panchromatic sharpening enhances HSIs by injecting detailed spatial information from a high-resolution panchromatic (PAN) image. Conversely, multispectral and HSI fusion leverages multispectral images to help hyperspectral maps acquire spatial details. For instance, Shi et al. [76] utilized both multispectral and hyperspectral images as conditional inputs to a diffusion model, subsequently generating high-resolution hyperspectral maps by integrating information from these two image modalities. Similarly, Qu et al. [77] developed a conditional recurrent diffusion framework. This framework enables the spatial super-resolution process to be guided by the multispectral image and the spectral super-resolution process by the hyperspectral map, leading to the generation of high-resolution hyperspectral maps that boast complementary spatial and spectral information.
More contemporary research is increasingly focused on utilizing Panchromatic (PAN) images for HSI super-resolution [78, 79, 80, 81, 82, 83, 84]. For instance, Meng et al. [79] introduced a modal calibration module to enhance and extract features from both image types, subsequently using these enhanced features as conditional inputs for a diffusion model. Conversely, Cao et al. [78] posited that the unique information from different image modalities should not be mixed during processing. They instead designed two conditional modulation modules to separately extract coarse-grained style information and fine-grained frequency information, which are then used as conditional inputs.
Figure 8: Converged network fusion mode [82]
To fully leverage the unique information from different modalities, Li et al. [82] proposed a panchromatic sharpening network based on a biconditional diffusion model. In this architecture, both HSI and PAN images serve as independent conditional inputs, allowing the network to separately learn spectral features and spatial textures. Designed based on LDM, this network achieves excellent super-resolution (SR) performance while significantly reducing computational cost. Similarly, Rui et al. [84] also perform the sampling process in a low-dimensional space. Their method is rooted in the fact that a high-resolution hyperspectral map can be decomposed into the product of two low-rank tensors. First, one of these low-rank tensors is computed using the low-resolution hyperspectral map. This tensor, along with the LRHSI (low-resolution hyperspectral image) and PAN, is then conditioned and fed into a pre-trained diffusion model [85] to generate the other low-rank tensor. Unlike the aforementioned methods, this approach is a completely unsupervised deep learning method, meaning it doesn’t require high-resolution hyperspectral maps during any part of the process, making it highly feasible for practical applications.
4.2.2 Data Generation and Completion
In the realm of generation and complementation, while broad categories include text-to-image and image-to-image synthesis, the hyperspectral domain exclusively features image-to-image applications. Here, diffusion models leverage image bootstrapping to create new images. These models can be guided by various forms of input, such as maps [86] or semantic layouts [87, 88, 89, 90]. Even with only specific information in these guiding images, excellent performance in image generation is still observed. For instance, Espinosa and Crowley [86] demonstrated that training conditional diffusion models with maps and historical maps could generate more realistic satellite images. Similarly, Yuan et al. [87] successfully produced high-quality hyperspectral images guided by semantic masks, simultaneously addressing the inherent challenge of lengthy training times for diffusion model convergence.
However, generating high-quality images alone isn’t always sufficient; creating more practical image-annotation pairs is often crucial. To this end, Zhao et al. [89] implemented a two-stage fine-tuning of the Stable Diffusion (SD) model, enabling generation from noise to annotations and then from annotations to images. Furthermore, Toker et al. [90] extended the standard diffusion model into a joint probabilistic model for images and their corresponding labels, allowing for the simultaneous generation of both.
Hyperspectral images (HSIs) offer rich spectral and spatial information, but their acquisition cost is significantly higher than that of multispectral images. Consequently, scientists have explored using readily available multispectral remote sensing images to generate HSIs [91, 92]. However, a challenge arises because the diffusion models used for this task require the input noise dimensions to match the spectral bands of the hyperspectral map. This can lead to an excessively large noise sampling space, making model convergence difficult.
To address this, Zhang et al. [91] introduced a spectral folding technique that converts the input hyperspectral map into a pseudo-color image before training the diffusion model.Inspired by the LDM, Liu et al. [92] utilized a Conditional Vector Quantization Generative Adversarial Network (VQGAN) [93] to derive a latent space for the hyperspectral map within the Latent Diffusion Model. Diffusion model training and sampling are then performed within this more manageable latent space.Furthermore, to tackle the issues of high cost and low spatial resolution, Deng et al. proposed DiffUn [94], a Diffusion Model-based hyperspectral spectral unmixing method. This approach combines spectral prior distributions learned by an unconditional Diffusion Model with likelihoods derived from the Linear Mixing Model (LMM) to enhance deblurring performance and synthesize hyperspectral maps.More recently, Dong et al. introduced a novel ISPDiff [95] model. This transparent, unified, and scale-flexible model addresses these problems using an iterative optimization algorithm. It incorporates a Multi-Scale Residual Network (MSRN) for scale-matching recovery and a Cross-Modal UNet (CMUN) for cross-scale upsampling.
Overall, image-to-image generation methods based on conditional diffusion models [67, 68] fundamentally generate images from noise, rather than directly converting existing ones, by using bootstrap images as conditional inputs to conditional DDPMs. To achieve a true image-to-image conversion, Yu et al. proposed UnmixDiff [96]. This model explores the feasibility of depth generative models in HSI (Hyperspectral Image) synthesis, synthesizing HSIs from the perspective of material distribution and introducing a clear physical interpretation into the task. By integrating an unmixing autoencoder within the diffusion generation model, hyperspectral image synthesis is projected from the high-dimensional spectral domain to a low-dimensional abundance domain. This significantly reduces computational complexity and effectively boosts the spectral fidelity of the generated HSIs. Constructing a diffusion generation model in the abundance space also allows for the generation of HSIs with realistic spatial details while maintaining spectral consistency. This type of approach notably employs self-supervised learning to extract feature embeddings from input patches. These embeddings then serve as conditional inputs, guiding the diffusion model’s learning process. Crucially, this method preserves the spatial arrangement of each color patch from the original image, enabling the generated patches to be seamlessly assembled into a large and coherent image.
Figure 9: UnmixDiff generates a simple diagram of the HIS [96]
4.2.3 Classification
Among applications in the hyperspectral field, the classification task is paramount for image analysis, enabling accurate identification and characterization of diverse structures and anomalies within an image. It significantly aids professionals in interpreting large volumes of complex data. Despite this potential, image classification aims to extract more valuable information about detailed features, such as buildings or grass, by assigning each pixel to a specific class [97, 98]. However, hyperspectral images inherently contain highly diverse and complex scenes and information, which complicates accurate classification. Therefore, leveraging diffusion models to enhance classification remains a significant challenge that requires further strengthening and resolution. Nevertheless, because diffusion models can learn and simulate more intricate data distributions than other deep learning models, there’s a growing commitment to applying them to hyperspectral image classification, leading to increasingly promising results [99, 100, 101, 102, 103, 104].
For image classification, the complex data distributions inherent in hyperspectral images present greater challenges for traditional deep learning models. Fortunately, diffusion models excel at capturing the joint spectral-spatial features of these images [105, 106, 107], leading to improved classification accuracy [108, 109] and enhanced generalization ability [110].
Building on these strengths, Zhou et al. [111] utilized the temporal information of diffusion models to construct a time-stepping feature library. They also introduced a dynamic fusion module to combine spectral-spatial features with temporal features, allowing for sufficient image information before classification. Li et al. [112] designed a two-branched diffusion model that extracts features from both HSI and LiDAR images, achieving information complementarity across different modalities. This approach boosts pixel discriminability and demonstrates excellent performance in more complex hyperspectral tasks [37]. Qu et al. [113] further deepened this by configuring the encoder of their two-branch diffusion model to operate in a parameter-sharing mode, ensuring shared features are extracted from multimodal images. Additionally, Chen et al. [114] leveraged diffusion models to assist in deep subspace construction, achieving impressive unsupervised HSI classification performance.
Taking a different approach, Qu et al. [115] combined a diffusion model with both a super-resolution (SR) network and a classification network. This allowed for the progressive reconstruction of high-quality images by iteratively performing image super-resolution and classification tasks, creating a mutual guidance mechanism. By utilizing multi-scale classification results to guide the target-scale image in learning new knowledge and further reconstructing high-quality images, both super-resolution and classification performance were significantly improved. Jiang et al. proposed an image fusion classification network based on Denoising Diffusion Probabilistic Models (DDPMs) [107], which demonstrated a significant improvement in classification accuracy after fusion. Similarly, Zhu et al. introduced a Spatial-Spectral Diffusion Comparison Representation Network (DiffCRN), also based on DDPMs [102]. This network incorporates a spatially self-attentive denoising module (SSAD) and a spectral group self-attentive denoising module (SGSAD). It enhances the efficiency of spectral-spatial feature learning by designing new DDPM loss functions and Contrastive Learning (CL) with Log Absolute Error (LAE). Furthermore, the inclusion of an adaptive weighted addition module (AWAM) and a cross-time step spectral spatial fusion module (CTSSFM) effectively mitigates interference from spatial spectral heterogeneity and noise effects, leading to better classification.
Figure 10: Diffusion model based image classification process diagram
4.2.4 Segmentation
Image segmentation is a crucial task in computer vision, encompassing areas like semantic segmentation [116] and general image segmentation [14, 117]. In deep learning, semantic segmentation is considered a subtask of image segmentation [112, 118]. Hyperspectral image segmentation specifically aims to simplify image complexity by breaking it down into multiple meaningful segments. This process provides valuable information for hyperspectral-related fields, thereby facilitating image analysis. However, deep learning models typically demand a large volume of diverse pixel-annotated training data to yield generalizable results. Unfortunately, the number of images and labels available for hyperspectral image segmentation is limited due to the significant time, cost, and expertise required for annotation. Consequently, diffusion modeling has emerged as a promising avenue in image segmentation research, as it can synthesize labeled data, thereby eliminating the need for extensive pixel annotation [102, 104].
Dong et al. [95] are utilizing Denoising Diffusion Probabilistic Models in conjunction with Scale Matching Recovery Networks (P2MSRN) and cross-scale model-driven up-sampling networks. Chen et al. [119] are designing image conditional mixing guides and integrating them with Denoising Diffusion Probabilistic Models. For labeling information, Zhang et al. [110] leverage Diffusion Probabilistic Models, while Zhu et al. [120] employ Variational Auto-Encoders alongside semi-supervised diffusion models for annotation. Li et al. proposed a multidomain diffusion-driven feature learning network (MDFL) [112] to enhance pixel-level classification (i.e., segmentation) performance of high-dimensional remote sensing images through a multidomain diffusion model and feature reuse mechanism. Zhu et al. also proposed DiffCRNs [102], a diffusion-based model coupled with contrastive learning, to extract spatial-spectral features and ultimately generate pixel-level classification circles. Furthermore, Xu et al. introduced a data-enhanced semantic segmentation framework based on the diffusion model [104]. This framework generates pseudo-labeled data via the diffusion model, uses ControlNet for cross-modal "label→multispectral image" generation, and employs multi-stage training to improve segmentation accuracy.
4.2.5 Target Detection
In contrast to pixel-level ground classification, target detection for hyperspectral images focuses on localizing specific instances, such as an airplane, vehicle, or ship, within the image using bounding boxes [121, 122]. A significant challenge in this area is the ambiguity and lack of realism and diversity in training data, often exacerbated by external disturbances [123, 118]. This issue stems from the inherent nature of hyperspectral images, where objects needing detection are frequently complex, difficult to distinguish, and distributed across various image regions. Consequently, augmenting the objects to be detected using diffusion models has emerged as an effective solution.
To address the aforementioned challenges, Yu et al. proposed Diff-Mosaic (Enhanced Target Detection via Diffusion Prior) [118]. This diffusion model-based approach introduces an enhancement network called Pixel-Prior, which coordinates pixels to generate highly realistic and diverse mosaic images. It leverages a powerful diffusion model prior that resamples pixels with samples, thereby introducing real-world information and ensuring the generated samples are both realistic and varied.
Similarly, Qi et al. presented a novel joint framework for underwater hyperspectral image restoration and target detection, based on a conditional diffusion framework [123]. This framework employs a variable spectral group extraction module and a joint underwater hyperspectral image restoration and target detection (JURTD) module to simulate different ranges of underwater target spectra. This achieves high-quality restored images and superior detection performance, effectively overcoming the inherent challenges of underwater hyperspectral imaging. Zhu et al. introduced ODGEN (Domain-specific Object Detection Data Generation with Diffusion Models) [124], a novel approach to generate high-quality, bounding-box-conditioned images. This facilitates data synthesis specifically for object detection. By utilizing synthetic visual cues with spatial constraints and textual descriptions per object to control the diffusion model, ODGEN demonstrates robust performance when handling complex scenes and specific domains.
Meanwhile, to enhance precision and accuracy, Wang et al. proposed OrientedDiffDet (pairwise object-oriented diffusion model) [125]. This approach combines oriented frame representation, feature alignment, and accelerated diffusion, allowing for direct object detection from a set of random frames. During the training phase, horizontal detection boxes are initially converted to oriented detection boxes. The model then learns to reverse this conversion by diffusing from ground truth oriented boxes to a random distribution. In the inference phase, the model progressively refines a set of randomly generated boxes to produce the final output.
Figure 11: OrientedDiffDet target detection diagram with targets boxed in different colors in the detection diagram [125]
4.2.6 Anomaly Detection
Image anomaly detection, a critical area in computer vision, focuses on highlighting unusual regions within an image [115, 126, 127, 128]. In recent years, generative modeling has significantly influenced anomaly detection research, showing promising results. This section will explore how diffusion models, as a prominent generative model, can be applied to hyperspectral image anomaly detection. While studies specifically on diffusion model-based hyperspectral image anomaly detection are relatively few, there are numerous applications in other fields, such as medical imaging [129, 130] and industrial inspection [131, 132]. Given the ongoing advancements in diffusion modeling, its use in hyperspectral image anomaly detection is expected to grow significantly in the future.
In the field of anomaly detection, a persistent challenge is the complexity of background clutter in hyperspectral images, which often obscures anomalies. Additionally, the scarcity of labeled hyperspectral samples leads to poor generalization in existing anomaly detection methods.
To tackle these issues, Ma et al. proposedBSDM (Background Suppression Diffusion Model) [133], a novel anomaly detection solution. BSDM simultaneously learns the latent background distribution and generalizes it across different datasets to suppress complex backgrounds. This approach is characterized by three key aspects: 1. Pseudo background noise is designed for the complex backgrounds of HSIs, and a Diffusion Model (DM) is employed to learn the underlying latent background distribution. 2. To address generalizability, a statistical bias module adapts BSDM to datasets from various domains without requiring labeled samples. 3. For effective background suppression, the DM’s inference process was innovatively improved by feeding the original hyperspectral image into a denoising network, which removes the background as if it were noise. This work also marked the first application of diffusion modeling to hyperspectral image anomaly detection.
Wu et al. introduced the diffusion background dictionary method (DBD) [134], combining the strengths of model-driven and data-driven approaches based on HSI and HAD (Hyperspectral Anomaly Detection) task characteristics. DBD intrinsically integrates the DM with the low-rank representation (LRR) model, utilizing the DM to obtain the crucial background dictionary tensor within the tensor LRR, thus enabling accurate anomaly detection. Following the principles of the RX algorithm in anomaly detection, the multivariate normal distribution approximating the HSI background is diffused to better suit background suppression. Furthermore, hyperspectral image anomaly detection often grapples with a lack of prior information, primarily the challenge of accurately reconstructing background regions while inferring the latent background of anomalous areas. To this end, Chen et al. proposed Double Window Spectral Diffusion (DWSDiff) [56]. This novel approach models accurate background estimation through an iterative spectral diffusion and reverse diffusion process. A dual-window strategy is also introduced to mitigate the impact of extended anomalous regions in the neighborhood on background estimation. Finally, an anomaly generation strategy based on Principal Component Analysis (PCA) and the Linear Spectral Mixture Model (LSMM) is devised to create realistic training data, overcoming the scarcity of labeled hyperspectral data and leading to improved performance.
Figure 12: Various diagrams under BSDM from left to right: anomaly diagram, 1D t-SN diagram, 2D t-SNE scatter diagram [133]
In conclusion, the application of diffusion models in hyperspectral image anomaly detection remains a rapidly evolving field with significant research potential. We anticipate a growing number of researchers will focus on this area, driving substantial advancements in anomaly detection.
4.2.7 Noise Suppression
A primary challenge in hyperspectral imaging is acquiring images without losing crucial information. During acquisition or subsequent processing, images can be corrupted by noise or artifacts, significantly degrading quality, especially with small objects or low contrast. Due to their generative nature, diffusion models are well-suited to addressing various noise reduction problems [135, 136, 137]. For instance, [136] focuses on denoising in image restoration by designing a multi-shape spatial rectangular self-attention module and a spectral latent diffusion enhancement module. Leveraging the diffusion model allows for a tighter alignment with the desired HSI, enhancing denoising robustness.
In image processing, hyperspectral images (HSIs) are invariably affected by various types of noise due to the inherent limitations of imaging techniques and environmental factors. Fundamentally, the learning process of diffusion models is akin to a denoising process [19], making them exceptionally effective for noise reduction in hyperspectral contexts. Different image modalities face unique noise challenges. For instance, HSI recovery can be particularly susceptible to distributional biases and often suffers from data scarcity [138]. HSIs also contend with the inhomogeneous distribution of noise along the spectral dimension [139] and the correlation of noise across different frequency bands [140, 141]. Fortunately, diffusion models can efficiently address these noise issues.
Ceng and Cao et al. proposed Diff-Unmix [142] to tackle the challenge of recovering noise-degraded HSIs by fusing spectral unmixing with conditional abundance generation. This method employs a learnable block-based spectral unmixing strategy, supported by a purely transformer-based backbone. They then introduce a self-supervised generative diffusion network to augment the abundance maps from the spectrally unmixed blocks, reconstructing noise-free decomposition probability distributions and effectively mitigating noise-induced degradation within these components. Finally, the HSI is reconstructed through de-mixing by combining the diffusion-adjusted abundance maps with the spectral end-members. He et al. [140] introduced a truncated diffusion model that initiates denoising from an intermediate step of the diffusion process, rather than from pure noise. This approach aims to preserve the inherently valid information within the HSI. Additionally, Deng et al. proposed RGB image-enhanced hyperspectral Denoising Diffusion Probabilistic Models [143], improving DM-based HSI denoising by integrating RGB-DM pre-training on supplementary RGB images. They also introduced a fusion operator to combine HSI-DM and RGB-DM, further enhancing the performance of DM-based HSI denoising. Kai et al. proposed a multi-scale spatial spectral joint denoising network for hyperspectral images (MSFDN) [137]. To more accurately describe and extract noise for superior denoising, they constructed a subspace projection network to convert HSIs into a low-dimensional subspace. They also designed a noise separation network and a subspace inverse projection network, leveraging a conditional diffusion model for enhanced robustness.
Figure 13: Process diagram of noise suppression based on diffusion model
4.2.8 Data Recovery
Hyperspectral Image (HSI) recovery aims to reconstruct clean images from degraded observations [144], playing a pivotal role in subsequent HSI applications. This recovery process faces significant hurdles in effectively leveraging the intrinsic spatial nonlocal self-similarity and spectral low-rank properties of HSIs, further complicated by existing data scarcity. Traditional model-based approaches struggle to accurately model complex image features with manual prior assumptions, while deep learning methods often suffer from poor generalization, model complexity, suboptimal results, and parameter redundancy. However, diffusion models have recently garnered considerable attention for their impressive image recovery performance [145, 146, 147, 148, 149, 150, 151], particularly for their robustness to noise [92, 95, 140, 152, 115, 119]. It’s worth noting, though, that diffusion methods are typically trained on large datasets and perform exceptionally well within those distributions, but they can be quite susceptible to distributional bias.
To tackle these issues, Miao et al. introduced DDS2M (Denoising Diffusion Probabilistic Models for Spatial Spectroscopy) [138], a self-supervised diffusion model for HSI recovery. This model operates by inferring the parameters of its proposed variational spatial spectral module (VS2M) during reverse diffusion, using only the degraded HSI without needing any additional training data. A custom loss function based on variational inference enables untrained spatial and spectral networks to learn a posteriori distributions, which serve as transitions in the sampling chain to aid the reverse diffusion process. Chen et al. [56] employ a dual window, iterative diffusion model, and reverse reconstruction to mitigate the impact of background estimation. Li et al. proposed HIR-DIFF (HSI Recovery via Diffusion Model with Untrained Prior) [141], a supervised HSI recovery framework. This framework, based on singular value decomposition (SVD) and rank-revealed QR (RRQR) decomposition, recovers clean HSIs from the product of two low-rank components. They also introduced a novel exponential noise schedule to accelerate the recovery process. Furthermore, Li et al. presented a spectral latent diffusion enhancement module [136], which incorporates a latent diffusion-enhanced rectangular transformer for HSI recovery. Their proposed multi-shape spatial rectangular self-attentive module leads to a tighter alignment with the desired HSI.
Meanwhile, in the realm of data recovery, problems often arise from noise mixing, characterized by both signal-correlated and signal-uncorrelated components, which can affect visual perception and cause discriminative image distortion. To address this, Si et al. proposed CASSIDiff [153] for reconstructing CASSI metrics. This model employs a feature fusion mechanism based on the Discrete Wavelet Transform (DWT) and a spatial-spectral attention mechanism, demonstrating high reliability in recovery tasks. Yi et al. introduced an efficient diffusion model (EDiffSR) [149] for RSI (Remote Sensing Image) Super-Resolution. They developed an efficient activation network (EANet) and a practical conditional prior enhancement module (CPEM), which collectively reduce the computational budget while retaining more information. Meng et al. proposed FastDiffSR [148], a conditional diffusion model-based SRSISR (Single-Image Super-Resolution) method. FastDiffSR significantly improves generation quality and visual clarity by reducing the number of sampling steps, utilizing residual images, and combining channel and spatial attention mechanisms.
Figure 14: Diffusion model-based research progress in hyperspectral image processing and analysis the tasks are established in nine subfields: (1) data enhancement and synthesis, (2) data generation and complementation, (3) classification, (4) segmentation, (5) anomaly detection, (6) target detection, (7) noise suppression, and (8) data recovery. No abbreviated name, with two initial letters at the beginning and end
5 Challenges and Limitations of the Diffusion Model
5.1 Computational complexity and Time cost of the model
Diffusion Models achieve data generation and denoising through multi-step iterative Markov chains, whose computational complexity increases exponentially with the data dimensions. Taking the denoising diffusion probabilistic model (DDPM) [19] as an example, hundreds of forward and reverse diffusion computations are required to generate a single hyperspectral image, especially when processing hyperspectral data above 200 bands, and the GPU memory occupancy can be more than 20 GB (e.g., 1024 × 1024 resolution scenes). Although the improved DDIM [35] reduces the number of sampling steps to less than 50 through non-Markov chains, the spatial-spectral three-dimensional nature of hyperspectral data (see Section 2.2) still leads to a single inference time as long as several minutes, which makes it difficult to satisfy real-time demanding tasks, such as agricultural monitoring and disaster warning. Although the Latent Diffusion Model (LDM) [18] reduces the computational volume by compressing the latent space, the information compression in the spectral dimension may lose key features (e.g., mineral absorption peaks), which affects the classification and anomaly detection accuracy (see Table 2).
5.2 Impact of high dimensionality of hyperspectral data on model training
The performance of diffusion models is highly dependent on large-scale labeled datasets, and the cost of labeling hyperspectral data far exceeds that of traditional RGB images. For example, field sampling is required to validate crop classes in agricultural applications, and the cost of labeling a single image can be hundreds of dollars. Although DDPM [19] generates synthetic samples through data augmentation, its fidelity is limited by the distribution of training data, and it is prone to Mode Collapse in scenarios with fewer samples (e.g., rare mineral detection). Experiments show that when the training samples are less than 1000, the classification accuracy of the diffusion model decreases by about 12% (F1-score from 0.85 to 0.75) compared to Variational Auto-Encoder (VAE) [9]. In addition, conditional diffusion models based on classifier guidance (e.g., ADM-G [11]) need to pre-train high-precision classifiers, but the class imbalance problem of hyperspectral data (e.g., less than 1% of anomaly samples in anomaly detection) leads to difficulties in classifier training, which restricts the model generalization ability.
5.3 Adaptability issues to large-scale data
Real hyperspectral images often contain non-Gaussian noise (e.g., sensor thermal noise, atmospheric scattering noise), and most of the existing diffusion models are based on additive Gaussian noise assumptions (see Section 3.1.1) The forward process design of the DDPM [19] is only applicable to isotropic Gaussian noise, and the denoising effect is significantly degraded for streak or PoissonGaussian hybrid noise (PSNR reduction of 8 15 dB). Although Score-SDE [21] extends the scope of noise modeling through stochastic differential equations, its inverse process relies on the accurate estimation of the score function, which is susceptible to error accumulation in bands where the signal-to-noise ratio (SNR) fluctuates drastically (e.g., near-infrared bands). For example, the Diffusion Models’ denoising performance degrades by about 20% compared to the conventional wavelet approach under extreme conditions where the noise standard deviation exceeds 30 dB. In addition, the mixed image element problem (see Section 2.3) leads to a high degree of spatial coupling between the noise and the target signal, further exacerbating the modeling difficulty.
5.4 Limitations of current models in practical applications
The black-box nature of diffusion models exposes them to credibility challenges in hyperspectral critical applications. For example, in anomaly detection tasks, the model may misclassify rare features as noise by overfitting the training data (literature [16]). Experiments have shown that LDM [18] based anomaly detectors have a 18% higher false alarm rate than the traditional RX algorithm (see Section 2.3) in the detection of unknown mineral types. In addition, although the Latent Diffusion Model improves the computational efficiency, the physical meaning of the latent vectors is unclear, which makes it difficult for users to trace the generation logic of the detection results. In scenarios that require strict validation, such as military target identification, the interpretability disadvantage of diffusion models significantly constrains their practical application. Comparative studies have shown that traditional methods (e.g., SVM) [154] can provide a decision basis through feature importance analysis, whereas diffusion models can only output probability estimates and lack a transparent inference chain.
In summary, the practical deployment of diffusion models in hyperspectral image processing is still limited by core bottlenecks such as computational efficiency, data dependency, noise robustness and interpretability. Future research needs to combine hyperspectral data characteristics (e.g., spectral continuity, spatial-spectral correlation), develop lightweight architectures and adaptive noise modeling methods, and explore semi-supervised learning frameworks to reduce annotation costs, so as to promote its large-scale application in remote sensing, agriculture and other fields.
6 Comparative and Experimental analysis
In order to effectively illustrate the superiority of the diffusion model in processing hyperspectral images, this paper takes the experimental results of classification, target detection, anomaly detection, image recovery, and noise suppression as an example, and evaluates the performance of the diffusion model and other existing techniques through the visualization results and quantitative indexes.
Table 2: Comparison of diffusion and traditional deep learning models for HSI
Deep Learning Model Advantages Disadvantages Future Development Trends
Diffusion Model Generates high-quality image details. Specializes in denoising and data reconstruction. Suitable for dealing with high-dimensional features of hyperspectral images. High computational cost. Slow training and inference. Sensitive to hyperparameters. Optimizing computational efficiency. Combining physical models to improve reconstruction accuracy. Lightweight design to adapt to real-time processing needs.
Generative Adversarial Network (GAN) Generate realistic hyperspectral data to support data augmentation. Enhance robustness by adversarial training in classification tasks. Training is unstable (pattern crashes). Generating data may have artifacts. Improved training stability (e.g., Wasserstein GAN). Combined with attention mechanisms to enhance local feature generation.
Variational Auto-Encoder (VAE) Effective degradation and feature extraction. Generating latent space representations. Supporting unsupervised learning. Generated images are fuzzy and do not retain enough detail. limited ability to model complex distributions. Improved decoder structure for enhanced detail generation. Combined with diffusion model to improve generation quality.
Autoregressive Models (AR) Strong sequence modeling capability for pixel-by-pixel prediction. Excellent performance in temporal hyperspectral analysis. High computational complexity (cannot be parallelized). Difficult to handle large size hyperspectral images. Introducing sparse attention mechanism to reduce computation. Combining with Transformer architecture to improve the efficiency of long sequence modeling.
6.1 Experimental results and Performance assessment and Various metrics
6.1.1 Introduction to evaluation metrics and Datasets
The receiver operating characteristic curve (ROC) is a qualitative analysis of the model. The horizontal coordinate of the curve is the false positive rate (FPR) and the vertical coordinate is the true positive rate (TPR).
AUC (area under ROC curve) is the area under ROC curve, which is a quantitative analysis of the model. In practice, AUC is often used as a metrics for evaluating the accuracy of a method. The larger the AUC value, the better the detection performance of the algorithm, with mAUC representing the average AUC. where larger ${\rm AUC}_{(D,F)}$and ${\rm AUC}_{(D,\tau)}$ indicate better detection performance, and are combined together defined as ${\rm AUC}_{TD}$; ${\rm AUC}_{BS}$ represents the ability of the detector to minimize false alarms due to background interference; ${\rm AUC}_{ODP}$ denotes the overall probability of detection; ${\rm AUC}_{TD-BS}$ specifically evaluates the performance of the detector in recognizing a targets in background noise, taking into account both detectability and background suppression; ${\rm AUC}_{SNPR}$ reflects the detector’s ability to distinguish between signal and noise.
OA denotes the ratio of the number of correctly classified samples to the total number of samples in all test samples; AA denotes the average of the classification accuracy for each category; k × 100 is the Kappa coefficient is a measure of the consistency of the classifier, which takes into account stochastic consistency and provides a measure of the consistency between the classifier and the ground truth.
IoU is a measure of the degree of overlap between the predicted frame and the real frame; $\mathbf{P}_{d}$ is defined as the ratio of the number of correctly detected targets to the total number of targets; and $\mathbf{F}_{a}$ is defined as the ratio of the number of false alarms to the total number of pixels in the image.
PSNR is the Peak Signal to Noise Ratio, a measure of the difference in quality between a super-resolution processed image and the original high-resolution image.FID (Fréchet Inception Distance) is a metrics used to assess the quality of the generated images, measuring their similarity to each other. Patch Similarity evaluates the perceptual difference between the generated image and the original high-resolution image. parmas is the number of parameters for model training. entropy is usually used to measure the uncertainty and complexity of image information. niqe is the Natural Image Quality Evaluator, which evaluates the image quality metrics. SSIM (Structural Similarity) measures the similarity between two images, comparing images in terms of brightness, contrast and structure; FSIM (Feature Similarity) evaluates the image quality by calculating the phase coherence and gradient similarity in localized regions of the image, and is an improved version of SSIM; SAM (Sum of Absolute Errors) is another measure of the effectiveness of image denoising, which calculates the sum of the absolute value of the pixel differences between a denoised image and the original image counterparts.
Table 3: Description of benchmark datasets used for evaluation
Datasets Sensor Spectral Band Size Resolution
Salinas AVIRIS 204 512$\times$217 3.7 m
Pavia University ROSIS 103 610$\times$340 1.3 m
Indian Pines AVIRIS 200 145$\times$145 20 m
SIRST FLIR LWIR 512$\times$512 0.5 m
NUDT-SIRST Infrared Sensor LWIR 640$\times$512 0.3 m
Airport-Beach-Urban AVIRIS 205/188/191 150$\times$150 20 m
Potsdam ISPRS RGB(3) 5000$\times$5000 5 cm
Toronto Toronto RGB(3) 1920$\times$1080 0.2 m
UC Merced USGS RGB(3) 256$\times$256 0.3 m
CAVE Hyperspectral Camera 31 512$\times$512 N/A
CAVE-Toy Hyperspectral Camera 31 512$\times$512 N/A
KAIST Thermal Camera LWIR 640$\times$480 0.5 m
6.1.2 Experimental results
In this section, a selection of typical methods from those previously discussed will be chosen for a comparative study of their detection performance. This comparison will encompass traditional methods such as GANs, VAEs, autoregressive models, other deep learning approaches, and diffusion models. To ensure a more reasonable assessment of anomaly detection performance across different algorithms, a common dataset will be used for comparison whenever possible. In the results, the bolded values will indicate the optimal performance achieved by an algorithm on that specific dataset.
As illustrated in Figure 15, the classification performance of the first three methods is notably poor, exhibiting significant color adulteration across different classes. In contrast, the latter three methods demonstrate superior overall classification effects. Among these, the diffusion model-based method, SpectralDiff, achieves the best classification results, closely followed by DMVL. Both SpectralDiff and DMVL show improved boundary delineation compared to other methods, with minimal color adulteration. Although SpectralDiff has some minor classification issues in tiny areas, its overall performance is significantly better than other mainstream approaches.
Figure 15: Classification results on multiple datasets using different methods: (a) Pseudo-color image (b) Real labeled image (c) SF [155] (d) miniGCN [156] (e) SSFTT [157] (f) DMVL [158] (g) SSGRN [159] (h) SpectralDiff [108]
Table 4: Comparison of Metrics for Different Classification Methods on Indian Pines Dataset
Methods Evaluation metrics (Indian Pines)
OA AA K×$\times$100 Time(s)
SF 67.782 80.639 63.820 29.81
miniGCN 63.916 69.091 59.439 0.95
SSFTT 91.566 95.661 90.392 31.58
DMVL 90.487 81.311 89.216
SSGRN 92.336 96.099 91.254 1.13
SpectralDiff 93.146 96.438 92.175 48.53
Table 4 reveals that the Spectral-Diff method achieves the top position in the Indian Pines dataset classification across overall accuracy (OA), average accuracy (AA), and kappa coefficient. Its performance shows a significant improvement over other methods, marking it as the current optimal approach. Notably, its OA and AA are 0.8% and 0.34% higher than the second-place method (SSGRN), indicating superior robustness in classification, more effective feature capture, and reduced noise interference. However, a drawback of Spectral-Diff is its considerably higher time consumption compared to other methods.
Table 5: Comparison of Metrics for Different Classification Methods on Pavia University Dataset
Methods Evaluation metrics (Pavia University)
OA AA K×$\times$100 Time(s)
SF 81.511 84.942 75.795 163.34
miniGCN 82.355 84.037 76.620 2.40
SSFTT 93.667 93.448 91.617 110.08
DMVL 89.768 86.116 86.580
SSGRN 93.850 94.436 91.919 4.99
SpectralDiff 94.775 93.843 93.061 256.23
As Table 5 shows, this method achieves optimal results on the Pavia University dataset for two metrics, surpassing the second-place method (SSGRN) by 0.98% and 1.23%. While another metric is only 0.63% lower than the optimum, which is a relatively small difference, it remains the best among these methods. However, a significant drawback is its time consumption, which is more than double that of other approaches.
Table 6: Comparison of Metrics for Different Classification Methods on Salinas Dataset
Methods Evaluation metrics (Salinas)
OA AA K×$\times$100 Time(s)
SF 88.248 93.263 86.973 141.30
miniGCN 88.181 94.279 86.823 2.13
SSFTT 95.789 98.272 95.322 55.96
DMVL 97.005 95.853 96.668
SSGRN 96.539 96.354 96.144 2.98
SpectralDiff 98.971 99.465 98.854 242.16
Table 6 presents the classification results on the Salinas dataset, leading to the same conclusion as Tables 4 and 5. Across all three datasets (Indian Pines, Pavia University, and Salinas), the Spectral-Diff method consistently demonstrates optimal performance when compared to other approaches. However, this superior accuracy comes at the cost of higher computational time and slower processing speeds. If the goal is to balance accuracy with efficiency, SSGRN emerges as the optimal choice. Conversely, for applications demanding low latency and high tolerance for accuracy, miniGCN is the most suitable option.
As seen in Figure 16, the overall performance of these mainstream target detection methods is good. However, some methods exhibit instances of undetected points or false positives. Among them, the diffusion model-based method, Diff-Mosaic, demonstrates the best detection effect, providing the most accurate and detailed boundary descriptions for detected points. It is largely consistent with the true labeling map, with virtually no missing components.
Figure 16: Target detection results on various datasets with competing methods: (a) Original image (b) Real labeled image (c) ALC-Net [160] (d) ACM [161] (e) UIU-Net [162] (f) DNA-Net [163] (g) Diff-Mosaic [118]
Table 7: Comparison of Metrics for Different Target Detection Methods on SIRST Dataset
Methods Evaluation metrics (SIRST)
IoU $P_{d}$ $F_{a}$
ALC-Net 70.00 71.71 23.91
ACM 76.17 86.31 16.07
UIU-Net 72.03 98.10 26.15
DNA-Net 76.97 95.41 3.54
Diff-Mosaic 79.44 99.99 3.19
Table 7 demonstrates that the Diff-Mosaic method achieves optimal performance for target detection on the SIRST dataset. It surpasses the second-place methods (DNA-Net and UIU-Net) with a 3.11% higher intersection ratio and a 1.89% higher detection probability. Furthermore, its false alarm rate is 87.8% lower than UIU-Net. This indicates that Diff-Mosaic’s detection rate approaches 100% with minimal errors, making it the superior method among those compared.
Table 8: Comparison of Metrics for Different Target Detection Methods on NUDT-SIRST Dataset
Methods Evaluation metrics (NUDT-SIRST)
IoU $P_{d}$ $F_{a}$
ALC-Net 72.13 78.81 21.22
ACM 71.11 85.13 22.51
UIU-Net 89.00 98.73 6.02
DNA-Net 88.38 97.99 4.04
Diff-Mosaic 91.18 99.47 1.91
Table 8 further corroborates the findings from Table 7, showing that Diff-Mosaic consistently delivers optimal performance on the manually synthesized NUDT-SIRST dataset. This reinforces the method’s superior effectiveness among the compared approaches.
Table 9: Comparison of Parameters and Reasoning for Different Target Detection Methods
Methods Calculate the number of parameters Time (s)
ALC-Net 0.38 40.93
ACM 0.29 18.53
UIU-Net 4.7 43.42
DNA-Net 50.54 33.98
Diff-Mosaic 4.7 43.42
Table 9 indicates that while Diff-Mosaic has a comparatively lower number of parameters, it isn’t the most parameter-efficient method. More significantly, its inference time is the slowest among those compared, being 57.3% higher than the fastest method (ACM). This highlights a key limitation of Diff-Mosaic in practical applications where speed is critical. When considering the collective results from Tables 7, 8, and 9, the optimal choice for target detection depends on specific priorities. If the goal is to achieve the absolute best performance, Diff-Mosaic stands out. However, if balancing strong performance with efficiency is paramount, DNA-Net proves to be the most suitable option. Conversely, for applications demanding high tolerance for minor accuracy variations and low latency, ACM is the ideal choice.
As presented in Figure 17, when comparing the results of anomaly detection experiments, the diffusion model-based methods, BSDM and DWSDiff, demonstrate superior detection performance compared to traditional mainstream approaches. They achieve a high detection rate and a low false detection rate, largely aligning with the true labeling map. In contrast, other mainstream detection methods exhibit poorer detection effects, characterized by lower detection rates (missing existing anomalies) and higher false detection rates.
Figure 17: Comparative anomaly detection performance across hyperspectral datasets: (a) Pseudo-color map (b) Real label map (c) GAED [164] (d) RGAE [165] (e) Auto-AD [166] (f) DFAN-HAD [167] (g) DirectNet [168] (h) BSDM [133] (i) TGFA [169] (j)DWSDiff [56]
Table 10: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{(p,r)}$ $mAUC_{(p,r)}$ $mAUC_{(p,r)}$
GAED 0.90756 0.15556 0.02072
RGAE 0.87615 0.14519 0.02401
Auto-Net 0.95028 0.15303 0.00627
DFAN-HAD 0.96552 0.45211 0.09603
DirectNet 0.95193 0.12799 0.00484
BSDM 0.94486 0.17342 0.03055
TGFA 0.92655 0.14559 0.00629
DWSDiff 0.97427 0.55329 0.01635
Table 11: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{(p,r)}$ $mAUC_{gs}$ $mAUC_{opp}$
GAED 1.06312 0.8868 1.04240
RGAE 1.02135 0.8521 0.99734
Auto-Net 1.10331 0.9440 1.09704
DFAN-HAD 1.41762 0.8694 1.32160
DirectNet 1.07993 0.9470 1.07508
BSDM 1.11828 0.9143 1.08773
TGFA 1.07214 0.9292 1.06585
DWSDiff 1.52756 0.9579 1.51121
Table 12: Comparison of metrics for each anomaly detection methods
Methods Evaluation metrics
$mAUC_{rp-gs}$ $mAUC_{svps}$
GAED 0.13484 13.79983
RGAE 0.12119 14.38906
Auto-Net 0.14676 52.05284
DFAN-HAD 0.35608 5.27753
DirectNet 0.12315 49.52139
BSDM 0.14287 8.40576
TGFA 0.13830 32.42845
DWSDiff 0.53694 67.20618
As Table 10, 11, 12 illustrates, which presents the accuracy of anomaly detection methods, BSDM and DWSDiff are diffusion model-based approaches. Among these, DWSDiff ranks first in all metrics except for integrated detection performance (${mAUC}_{TD}$), where it is second only to DFAN. While BSDM’s data isn’t as outstanding (ranking third in integrated detection performance), its effectiveness is still commendable.
Table 13: Comparison of Reasoning Time for Each Anomaly Detection Method
Methods Reasoning time (s)
Airport Beach Urban Average Time (s)
GAED 0.0300 0.0496 0.0257 0.0289
RGAE 0.0135 0.0251 0.0137 0.0141
Auto-Net 0.0020 0.0032 0.0020 0.0022
DFAN-HAD 0.0232 0.0448 0.0269 0.0264
DirectNet 0.0145 0.02839 0.0153 0.0145
BSDM 1.2306 1.2740 1.2509 1.2421
TGFA 1.6190 1.5505 1.5236 1.5392
DWSDUR 0.8016 0.8169 0.8192 0.8094
Table 13 indicates that both diffusion models, BSDM and DWSDiff, exhibit significantly longer and slower inference times across various medium dataset scenarios. Their average times are 99.8% (BSDM) and 99.7% (DWSDiff) higher than the shortest time achieved by Auto-Net, translating to hundreds of times slower than the optimal. This highlights a considerable challenge in terms of computational efficiency. Nonetheless, if the primary goal is to maximize anomaly detection performance, DWSDiff remains the optimal choice, with BSDM also being a strong contender.
As depicted in Figure 18, the overall image recovery effects are relatively modest across all methods. However, diffusion model-based approaches (DDPM, EDiffSR, FastDiffSR) demonstrate superior handling of details, resulting in a recovery quality that is notably closer to the high-resolution ground truth.
Figure 18: Visual comparison of hyperspectral image recovery methods: (a) High resolution map (b) HESNet [170] (c) TransENet [171] (d) NDSRGAN [172] (e) HAT [173] (f) DDPM [19] (g) EDiffSR [149] (h) FastDiffSR [148]
Table 14: Comparison of Metrics for Different Image Recovery Methods on Potsdam Dataset
Methods Evaluation metrics(Potsdam)
PSNR FID LPIPS Params
HESNet 32.932 44.192 0.3014 5.4330
TransENet 32.768 41.310 0.2878 37.459
NDSRGAN 30.529 51.209 0.2137 17.510
HAT 32.987 44.932 0.2962 25.821
DDPM 31.989 24.307 0.0997 26.166
EDuFSR 30.860 36.669 0.1104 26.790
FastDuFSR 33.130 22.832 0.1152 23.229
As evidenced in Table 14, the three methods based on diffusion models—DDPM, EDiffSR, and FastDiffSR—demonstrate unique advantages in key image recovery indicators. Notably, FastDiffSR exhibits the best overall performance in terms of PSNR and FID, significantly outperforming traditional methods. Compared to the second-best methods (HAT for PSNR and DDPM for FID), FastDiffSR shows substantial improvements: 0.43% for PSNR over HAT, and 6.07% and 13.5% for FID over DDPM. This indicates that FastDiffSR recovers images with superior detail fidelity and distributional consistency. However, a drawback for all diffusion model-based methods is their higher number of parameters compared to HESNet, specifically 79.4% higher for DDPM, 79.7% for EDiffSR, and 76.6% for FastDiffSR. This increased parameter count contributes to longer training times.
Table 15: Comparison of Metrics for Different Image Recovery Methods on Toronto Dataset
Methods Evaluation metrics (Toronto)
PSNR FID LPIPS Params
HESNet 25.413 118.791 0.3234 5.5070
TransENet 25.565 97.958 0.3112 37.533
NDSRGAN 23.237 97.922 0.2534 17.510
HAT 25.368 119.05 0.2957 25.821
DDPM 24.733 66.583 0.1759 26.166
EDdfSR 23.990 81.412 0.1626 26.790
EasDdfSR 25.527 55.761 0.1561 23.229
Similar to Table 14, Table 15 also demonstrates that the three methods based on the diffusion model exhibit strong overall performance, with FastDiffSR again standing out as the best. When combining insights from both tables, it’s clear that these three diffusion methods produce higher quality generations and retain more semantic information compared to traditional approaches. However, a consistent issue across all three diffusion methods remains their relatively larger number of parameters.
Table 16: Comparison of Measurement metrics for Different Image Recovery Methods
Methods Measurement metrics
Entropy NIQE
HESNet 7.0826 19.913
TransENet 7.0853 20.319
NDSRGAN 7.0528 19.214
HAT 7.1177 20.385
DDPM 7.1204 16.532
EDiIfSR 7.1838 17.641
EastDiffSR 7.1254 16.342
Table 16, which presents entropy and generation quality metrics, reinforces the findings from Tables 14 and 15. It demonstrates that the three diffusion methods—FastDiffSR, DDPM, and EDiffSR—retain more semantic information and achieve higher generation and recovery quality. Overall, FastDiffSR offers a balanced performance in terms of both recovery accuracy and efficiency. DDPM holds an advantage in perceptual quality, while EDiffSR can be considered a practical choice for moderate performance needs.
As illustrated in Figure 19, when comparing the noise map and the true noiseless map, the Diff-Unmix method demonstrates the best overall performance in noise suppression, followed by DDRM and then DDS2M. All of these are diffusion model-based approaches, and they achieve notably successful noise suppression. In contrast, traditional mainstream methods exhibit poorer effects, with LLxRGTV being the best among them.
Figure 19: Performance comparison of denoising methods on hyperspectral images: (a) Noise map (b) True noise-free map (c) LLxRGTV [174] (d) FGSLR [175] (e) LRTDCTV [176] (f) DIP [177] (g) LLnRPnP [178] (h) HLRTF [110] (i) DDRM [179] (j)DDS2M [138] (k)Diff-Unmix [142]
Table 17: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Noise: N(0,0,3))
PSNR SSIM FSIM SAM Time (s)
LLxRGTV 27.640 0.679 0.868 0.218 38
FGSLR 25.561 0.480 0.718 0.470 499
LRTDCTV 24.593 0.533 0.739 0.431 42
DIP 20.063 0.405 0.798 0.538 74
LLnRPhP 25.102 0.592 0.768 0.425 289
HLRTF 30.340 0.689 0.874 0.329 25
DDRM 27.810 0.786 0.893 0.387 23
DDS2M 30.078 0.666 0.834 0.355 318
Diff-Unmix 31.408 0.902 0.958 0.282 42
Table 18: Comparison of Metrics for Different Noise Suppression Methods on KAIST Dataset
Methods Evaluation metrics(KAIST Noise: N(0,0,2))
PSNR SSIM FSIM SAM Time (s)
LLaRGUX 31.152 0.802 0.917 0.205 38
FGSLR 30.126 0.737 0.878 0.262 249
LRTDCTV 25.952 0.658 0.816 0.406 43
DIP 24.181 0.608 0.825 0.475 72
LLnRPhP 28.664 0.748 0.861 0.379 240
HLRTF 33.011 0.808 0.925 0.275 23
DDRM 29.412 0.865 0.922 0.293 20
DDS2M 32.804 0.786 0.895 0.334 354
Diff-Unmix 33.059 0.964 0.940 0.116 37
In the experiments detailed in Tables 17 and 18, the last three methods—DDRM, DDS2M, and Diff-Unmix—are all based on the diffusion model. When evaluated on the KAIST dataset under two different noise conditions, these three methods demonstrate distinct strengths in noise suppression. Diff-Unmix consistently shows the most outstanding performance across key indices, achieving the highest PSNR, SSIM, and FSIM values. This highlights its significant advantages in noise suppression, structure preservation, and spectral fidelity compared to other methods. While the processing time for diffusion model-based methods like Diff-Unmix is slightly higher (with HLRTF being the fastest), their overall efficiency remains within an acceptable range.
Table 19: Comparison of Metrics for Different Noise Suppression Methods on CAVE Dataset
Methods Evaluation metrics (CAVE)
PSNR SSIM FSIM SAM Time (s)
LLxRGTX 28.743 0.712 0.829 0.121 40
FGSLR 24.474 0.346 0.554 0.511 1758
LRTDCTV 24.394 0.444 0.631 0.315 47
DIP 20.480 0.554 0.864 0.313 66
LLnRPhP 24.809 0.445 0.668 0.284 246
HLRTF 30.308 0.770 0.861 0.185 133
DDRM 30.521 0.754 0.873 0.196 16
DDS2M 30.308 0.724 0.863 0.251 319
Diff-Unmix 32.714 0.940 0.957 0.129 43
Table 19 indicates that the diffusion model-based methods generally outperform other approaches, with Diff-Unmix remaining the most prominent. This consistently demonstrates the overall superiority of diffusion models in noise suppression, delivering optimal performance and the best noise suppression effects.
Table 20: Comparison of Metrics for Different Noise Suppression Methods on CAVE-Toy Dataset
Methods Evaluation metrics (CAVE-Toy)
PSNR SSIM FSIM SAM Time (s)
LLARGUX 26.450 0.752 0.874 0.118 39
FGSLR 21.525 0.573 0.798 0.241 1149
LRTDCTV 25.214 0.665 0.783 0.228 45
DIP 22.092 0.594 0.895 0.191 67
LLnRPhP 23.088 0.541 0.717 0.207 240
HLRTF 27.308 0.730 0.941 0.161 41
DDRM 27.886 0.858 0.910 0.159 16
DDS2M 29.344 0.844 0.977 0.111 320
Diff-Unmix 28.046 0.945 0.993 0.156 43
Table 20 largely corroborates the findings from Tables 17, 18, and 19, confirming that diffusion model-based methods generally offer the best noise suppression performance. When considering all these tables together, the overall effect of diffusion models is highly acceptable, demonstrating a clear denoising effect that closely approaches noiseless images while retaining superior detail. This is true despite their processing time not being exceptionally fast.
6.2 Key discoveries and Summaries from the Experiments
As evidenced by the experimental comparisons in Section 5.2.2, diffusion models exhibit significant advantages in hyperspectral image processing tasks, while also revealing certain limitations.
In classification tasks, the SpectralDiff method, based on diffusion models, demonstrates superior accuracy metrics (OA, AA, and Kappa coefficients) across the Indian Pines, Pavia University, and Salinas datasets. For instance, its OA values were 93.15%, 94.78%, and 98.97% respectively, outperforming traditional deep learning methods like GANs and VAEs. This indicates its effectiveness in capturing complex spectral-spatial features. However, a significant drawback is its inference time, which is considerably higher (e.g., 48.53 s on Indian Pines) compared to lightweight models such as miniGCN (0.95 s), exposing a computational efficiency bottleneck for high-dimensional data processing.
For target detection, the diffusion-based Diff-Mosaic method achieves optimal performance on the SIRST and NUDT-SIRST datasets, with excellent IoU (79.44% on SIRST), high detection probability (99.99% on SIRST), and a low false alarm rate (3.19% on SIRST). It particularly excels at complex background suppression and precise target localization. Despite this, its parameter count (4.7M) and inference time (43.42 s) remain higher than traditional methods like ACM (0.29M parameters, 18.53 s inference time). This suggests that further lightweighting of diffusion models is needed to meet real-time demands in target detection.
In anomaly detection, DWSDiff comprehensively leads in mAUC metrics (e.g., ${mAUC}_{TD}$ of 0.97427), validating its ability to suppress complex backgrounds through its dual-window strategy and iterative diffusion. While BSDM’s performance is also respectable, DWSDiff stands out. Nevertheless, DWSDiff’s inference time (0.8094 s) is still substantially higher than highly efficient methods like Auto-AD (0.0022 s). Additionally, the model’s sensitivity to hyperparameters (e.g., noise scheduling) and its need for further optimization in generalization ability indicate that practical deployment of diffusion models in anomaly detection requires a careful balance between accuracy and efficiency.
For image recovery, diffusion model methods such as FastDiffSR, EDiffSR, and DDPM significantly outperform conventional methods in key metrics like PSNR and FID. For example, FastDiffSR achieved a PSNR of 33.13 dB and an FID of 22.83 on the Potsdam dataset, demonstrating superior detail fidelity and distributional consistency. Similarly, on the Toronto dataset, FastDiffSR maintained its advantage with a PSNR of 25.53 dB and an FID of 55.76, confirming its strength in semantic information retention. However, the high number of parameters in these diffusion models (e.g., FastDiffSR with 23.23M parameters) leads to a substantial increase in training time. This highlights that while diffusion models possess strong generative capabilities, their computational efficiency and model complexity require further optimization for practical deployment.
In noise suppression tasks, diffusion model-based methods, including Diff-Unmix, DDRM, and DDS2M, show significant advantages across metrics like PSNR, SSIM, and FSIM. On the KAIST dataset, Diff-Unmix achieved a PSNR of 31.41 dB, an SSIM of 0.902, and an FSIM of 0.958, underscoring its benefits in noise suppression, structure preservation, and spectral fidelity. Its superiority was further validated on the CAVE dataset with a PSNR of 32.71 dB and an SSIM of 0.940. Despite their effectiveness in removing noise and retaining details, the long inference time of diffusion models (e.g., 42 s for Diff-Unmix) restricts their application in scenarios demanding high real-time performance, suggesting a need for improved computational efficiency.
Some key points are also found.1. Accuracy advantage: diffusion models are significantly mainstream deep learning methods (e.g., GAN, VAE) in tasks such as classification, target detection, and anomaly detection, especially in complex background suppression, high-dimensional feature modeling, and noise robustness.2. Computational cost: the computational complexity and inference time of diffusion models are generally high, e.g., SpectralDiff The inference time on Salinas dataset reaches 242 seconds, which is far more than miniGCN (2.13 seconds), limiting its application in real-time scenarios (e.g., disaster monitoring).3. Model adaptation: diffusion models are sensitive to hyper-parameters (e.g., noise scheduling, number of diffusion steps), and need to be customized with optimization strategies for hyperspectral data characteristics. For example, DWSDiff mitigates the background estimation bias by a double-window strategy, but its generalization still needs to be verified.4. Complementarity with traditional methods: traditional methods (e.g., SVM) are still advantageous in terms of interpretability and lightness, while diffusion models are more suitable for high-precision and complex scene tasks. Future research can explore hybrid architectures to balance efficiency and performance.
7 Outlook
In recent years, diffusion models have made significant progress in tasks such as image generation, denoising and reconstruction by virtue of their powerful generative capabilities and modeling advantages over complex data distributions. The diffusion model shows unique potential in the multidimensional and high-information-density field of hyperspectral image processing. Hyperspectral images contain spectral information in hundreds of consecutive bands, but their high-dimensional characteristics and noise interference pose challenges for traditional methods. As shown in the previous chapter, diffusion models can accurately model spectral-spatial features through a step-by-step denoising process, thus achieving higher accuracy in the task. However, the fact is that diffusion models still have a long way to go in the hyperspectral domain, and there are many tasks that still need to be explored and further improved, and the current research is still limited by the consumption of computational resources, insufficient cross-modal fusion and lack of interpretability. In the future, technical innovations can be promoted in the following three aspects: innovation in fusion with other traditional deep learning models, optimization and deployment in real scenarios, and automated model design and hyperparameter tuning.
1, Fusion innovation with other deep learning models, so that the diffusion model and the mainstream architecture to form a complementary, such as CNN + diffusion model, the use of CNN to extract spatial-spectral features as the a priori conditions of the diffusion process, to improve the efficiency of the generation; Transformer + diffusion model, through the attention mechanism to model the long-range dependence, to optimize the super-resolution reconstruction of hyperspectral images; GAN / VAE + diffusion model, which first generates low-resolution samples with GAN or VAE, and then performs detail enhancement and denoising by diffusion model to balance the generation speed and quality; Physics-driven fusion: combining physical a priori such as radiative transfer model, to enhance the generalization ability of diffusion model in complex environments (e.g., atmospheric interference).
2. Optimization and deployment in real-world scenarios need to be achieved through model lightweighting (e.g., knowledge distillation, dynamic diffusion step size), hardware adaptation (parallelized FPGA/GPU-based design) and embedded deployment (low-precision inference on edge devices). For vertical domains such as agricultural disease monitoring and mineral exploration, task-oriented diffusion frameworks need to be developed to balance accuracy and real-time requirements.
3. Automated model design and hyperparameter tuning. The performance of diffusion models is highly dependent on noise scheduling, network architecture and training strategies. Automated techniques (e.g., neural architecture search, meta-learning) can adaptively optimize the model architecture; reinforcement learning-driven hyperparameter search can dynamically adjust the denoising process for different hyperspectral tasks. Meanwhile, the development of unified evaluation benchmarks (e.g., cross-dataset generalization metrics) will accelerate the application and iteration of automated techniques.
In conclusion, diffusion model provides a new paradigm for hyperspectral image processing, but its development still needs to break through the bottlenecks of computational efficiency, cross-modal fusion and domain adaptation. In the future, through the combination of algorithmic innovation, hardware synergy and automation technology, the diffusion model is expected to realize smarter and more efficient analysis systems in the fields of environmental monitoring, precision agriculture, etc., and to promote hyperspectral technology from the laboratory to large-scale applications.
8 Conclusion
In summary, diffusion models offer significant advantages over other deep generative models for hyperspectral image analysis. They demonstrate robustness to the inherent noise in hyperspectral images, adapt better to their variability and complexity, and provide a more stable training process. With their progressive denoising mechanism and high-dimensional data modeling capabilities, diffusion models have injected a revolutionary impetus into the field of hyperspectral image analysis. They effectively overcome the limitations of traditional methods in handling high-dimensional data, noise interference, and small sample scenarios through their Markov chain noise injection and inverse generation process. Therefore, the application of diffusion models in hyperspectral imaging is an inevitable trend. This paper first introduced relevant knowledge about hyperspectral images and the basic theory of diffusion models. Subsequently, it reviewed research on diffusion models within the hyperspectral domain, covering tasks such as data enhancement, generation and complementation, classification segmentation, anomaly detection, target detection, and denoising. Moreover, by comparing various visualization results and quantitative indices between diffusion models and traditional models in hyperspectral image processing, this paper fully demonstrated the superiority of diffusion models in this field. Finally, the limitations of existing diffusion models in the hyperspectral domain were discussed, emphasizing the need for further exploratory studies concerning extended applications and model deployment.